### v1

# Collaborative Unlabeled Data Optimization

Abstract
This paper pioneers a novel data-centric paradigm to maximize the utility of unlabeled data, tackling a critical question: How can we enhance the efficiency and sustainability of deep learning training by optimizing the data itself? We begin by identifying three key limitations in existing model-centric approaches, all rooted in a shared bottleneck: knowledge extracted from data is locked to model parameters, hindering its reusability and scalability. To this end, we propose CoOpt, a highly efficient, parallelized framework for collaborative unlabeled data optimization, thereby effectively encoding knowledge into the data itself. By distributing unlabeled data and leveraging publicly available task-agnostic models, CoOpt facilitates scalable, reusable, and sustainable training pipelines. Extensive experiments across diverse datasets and architectures demonstrate its efficacy and efficiency, achieving 13.6% and 6.8% improvements on Tiny-ImageNet and ImageNet-1K, respectively, with training speedups of 1.94×1.94$\times$1.94 × and 1.2×1.2$\times$1.2 ×  [^1] Our code is provided in the Supplementary Materials and will be publicly accessible..
1 Introduction
Figure 1: A Collaborative Data Optimization Framework CoOpt. For large-scale unlabeled data, self-supervised learning results in Refer to caption low training efficiency. Therefore, we propose CoOpt, an Refer to caption efficient and parallel framework enabling participants to use diverse task-agnostic models, such as pre-trained ResNets, termed prior models, for collaborative data optimization.
Deep Learning has achieved remarkable success, primarily due to the large-scale datasets [28, 35]. Despite the abundance of data in the era of big data, a significant portion of them remains unlabeled [22]. The dominant paradigm in the field for exploiting unlabeled data is self-supervised learning (SSL), which is fundamentally model-centric: it carefully crafted pretext tasks and loss functions to encode data information into model parameters [6, 13, 4].
However, the model-centric nature leads to three fundamental challenges. First, their training protocols are tightly coupled to specific network architectures, significantly hindering the transferability of trained model to other architectures [31, 18]. Second, even with advances in acceleration techniques, training over extensive unlabeled datasets still incurs prohibitively high computational costs [29]. Third, when a new downstream task arises, pretrained models typically require extensive fine-tuning or even full retraining. At the core of these challenges is a shared bottleneck: knowledge extracted from data is locked in model parameters, restricting its adaptability and preventing efficient reuse across diverse tasks or architectures.
To break free from model-centric paradigm, we propose a data-centric paradigm that directly optimizes the unlabeled data by optimizing targets for samples (further elaborated in Section 3.1), thereby effectively encoding knowledge into the data itself rather than into model parameters. The resulting “optimal data” is agnostic to downstream architectures, accelerates subsequent training by providing richer supervision, and can be reused across multiple tasks without repeated large-scale pretraining.
In the meanwhile, scaling this approach to massive unlabeled datasets introduces a significant challenge: a single node faces prohibitive compute and storage demands. To address this, we propose CoOpt, a highly efficient collaborative framework inspired by crowd-sourcing to achieve parallel data optimization. An overview of CoOpt is depicted in Figure 1, with detailed processes shown in Figure 2. In CoOpt, the unlabeled dataset is partitioned into disjoint subsets, each processed independently and in parallel by participants equipped with task-agnostic models. These models, referred to as prior models, can diverge from publicly available pre-trained models and the participants’ local models. Once the targets are optimized, they are aggregated to reconstruct a fully optimized dataset, achieving computational scalability through decentralized workload distribution.
CoOpt offers three key advantages: First, as the optimization relies solely on task-agnostic models, the aggregated optimized targets can be directly applied to any downstream architecture. Second, by distributing data subsets across participants, each node incurs only a fraction of the total computational load, enabling scalability and resource efficiency. Importantly, since only targets $D_{Y}$ are optimized, the aggregated targets are the same as those obtained by optimizing the whole dataset. Third, downstream tasks can directly fine-tune the optimized dataset for task-specific objectives without requiring training from scratch. Comprehensive experimental results have demonstrated that CoOpt achieves superior performance, generalization, and training efficiency.
In summary, our contributions are threefold:
  1. (a)
    We propose CoOpt, the first data-centric framework for optimizing unlabeled data, leveraging distributed computation to significantly lower the computational and storage costs per participant.
  2. (b)
    Within CoOpt, we identify a critical issue: Target Distribution Inconsistency in Section 3.3. Accordingly, we propose an effective target alignment strategy, elaborated in Section 3.4.
  3. (c)
    Extensive experiments have been conducted across a range of real-world scenarios, involving a variety of prior models trained on various datasets, architectures, and training paradigms. Notably, special cases are explored where human or significantly weak models are employed as prior models, verifying the robustness and flexibility of CoOpt. These experiments consistently demonstrate that CoOpt achieves superior effectiveness and efficiency.
2 Related Work
Model-centric Perspective: Self-supervised Learning.
It aims to exploit the intrinsic relationships within unlabeled data. For example, InstDisc [34] uses instance discrimination as a pretext task. MoCo [15] significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples. SimCLR [6] highlights the importance of hard positive sample strategies. Notably, BYOL [13] discards negative sampling and surpasses the performance of SimCLR [6].
Data-centric Perspective: Knowledge Distillation and Dataset Distillation.
Knowledge distillation [17] employs soft labels generated by teacher models to improve the performance of a student model and expedite its training [11]. Many following works [45, 44, 11] aim to enhance the use of soft labels. Dataset distillation [33] aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost. The majority of methods focus on optimizing images, which can be categorized into three primary approaches [22]: meta-learning frameworks [33, 46], matching-based methods [43, 42, 14] and decoupling frameworks [38, 29]. A more detailed description of knowledge distillation and dataset distillation is provided in Section A.
3 Collaborative Data Optimization Framework CoOpt
We begin by formally defining data optimization in Section 3.1. Subsequently, we provide a detailed description of the proposed CoOpt in Section 3.2. Furthermore, we identify an inherent challenge within this framework in Section 3.3 and present method in Section 3.4.
Table 1: Properties of Various Data Utilization Methods. ‘Optimize $D_{X}$’ and ‘Optimize $D_{Y}$’ indicate whether the optimization process encompasses the samples and targets, respectively. ‘SSL’ is self-supervised learning, ‘KD’ is knowledge distillation, and ‘DD’ is dataset distillation. ‘$\mathbf{c}$’ denotes the training cost of standard supervised learning that employs human-labeled data.
Method Optimize Efficiency Cost Analysis Total Efficiency
$D_{X}$ $D_{Y}$
Data
Optimization
Model
Training
Original SSL 0 [Uncaptioned image] ($\geq 2\mathbf{c}$) Extensive Computation [Uncaptioned image]
Labeled Data Optimization KD [Uncaptioned image] ($\geq\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Task-specific Teacher Models [Uncaptioned image]
DD [Uncaptioned image] ($\geq\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Task-specific Pre-trained models [Uncaptioned image]
CoOpt (Our) [Uncaptioned image] ($<\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Various Prior Models [Uncaptioned image]
3.1 Definition of Data Optimization
We first revisit current training acceleration techniques from a data-centric perspective, including knowledge distillation [17] and dataset distillation [33]. Specifically, we decouple data into samples $D_{X}$ and targets $D_{Y}$  [^2] A formal definition of $D_{X}$ and $D_{Y}$ is provided in Section 3.1., as summarized in Table 1. Essentially, compared to the self-supervised learning methods, these approaches achieve higher efficiency by either optimizing targets $D_{Y}$ pre-trained models (as in knowledge distillation), or by optimizing both input data $D_{X}$ and targets $D_{Y}$ (as in dataset distillation). Notably, for dataset distillation, optimizing input data $D_{X}$ is computationally more expensive than optimizing targets $D_{Y}$ [2]. Furthermore, recent studies [27, 25] have indicated that solely optimizing $D_{Y}$ not only reduces computational overhead but also achieves significant performance gains. These findings have demonstrated that optimizing $D_{Y}$ is both necessary and efficient. We refer to this process as data optimization.
Formally, consider a large-scale unlabeled dataset $D=D_{X}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{m}$ and $N=|D|$, data optimization aims to assign targets $D_{Y}=\{\mathbf{y}_{i}\}_{i=1}^{N}$ to construct an optimally labeled dataset $D^{\prime}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ such that models trained on $D^{\prime}$ can achieve higher performance of those trained on $D$ with significantly less training costs. This objective is expressed as
$\mathbb{E}_{(\mathbf{x},y)\sim P_{\mathcal{T}}}[\ell(\boldsymbol{\phi}_{\boldsymbol{\theta}_{D}}(\mathbf{x}),y)]>\mathbb{E}_{(\mathbf{x},y)\sim P_{\mathcal{T}}}[\ell(\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}(\mathbf{x}),y)]\,,$ (1)
where $P_{\mathcal{T}}$ denotes the test distribution, $\mathbf{x}$ is a test sample, $y$ is its label, $\ell$ is the loss function (e.g., cross-entropy loss), and $\boldsymbol{\theta}_{D}$ and $\boldsymbol{\theta}_{D^{\prime}}$ are parameters of network $\boldsymbol{\phi}$ trained on $D$ and $D^{\prime}$, respectively.
Notably, directly extending these methods to unlabeled data presents two challenges. First, they rely on task-specific models trained on labeled data, which are unavailable. Second, they incur substantial costs. For example, most dataset distillation methods rely on bi-level optimization [23]. To address these limitations, we propose a collaborative framework that leverages distributed computation and various task-agnostic models for unlabeled data. Specifically, inspired by [29], which shows that using task-agnostic models for target assignment can expedite training, we further enhance efficiency by splitting the data and then parallelly optimizing each split. After optimization, the optimal subset are aggregated to reconstruct a fully optimized dataset, achieving computational efficiency. Formally, we define data optimization with a prior model $\boldsymbol{\psi}$ in each participant in Proposition 1. {propositionframe}
Proposition 1 (Data optimization with prior model $\boldsymbol{\psi}$) .
Given samples $D_{X}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ and a prior model $\boldsymbol{\psi}:\mathbb{R}^{m}\to\mathbb{R}^{l}$, data optimization assigns optimal targets $D_{Y}=\{\mathbf{y}_{i}\}_{i=1}^{N}$ for the samples to create $D^{\prime}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$. We assigns a target $\mathbf{y}_{i}$ for $\mathbf{x}_{i}$ as:
$D^{\prime}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\mid\mathbf{y}_{i}=\mathbf{W}\boldsymbol{\psi}(\mathbf{x}_{i}),\forall\mathbf{x}_{i}\in D_{X}\}\,,$ (2)
where $\mathbf{y}_{i}$ is the optimized target, and $\boldsymbol{\psi}(\mathbf{x}_{i})$ represents the target of $\mathbf{x}_{i}$, which means the feature representation. $\mathbf{W}:\mathbb{R}^{l}\to\mathbb{R}^{n}$ denotes a matrix designed to transform the feature vector $\boldsymbol{\psi}(\mathbf{x}_{i})$ from dimension $l$ to $n$ without loss of information [24]. This transformation aligns the output dimension [^3] Here, $n$ is the target dimensionality of $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}$. In practice, each participant produces targets of varying dimensions due to using different prior models. Therefore, to train the model $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}$ on the optimized data $D^{\prime}$, we employ a random matrix $\mathbf{W}$ to transform all target vectors to a common dimensionality. with that required by the model trained on optimized data $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}:\mathbb{R}^{m}\to\mathbb{R}^{n}$.
Figure 2: Lifecycle of the proposed collaborative data optimization framework CoOpt. The framework encompasses an open data platform and multiple participants, involving five key data operations.
3.2 Overview of the Proposed Framework CoOpt
CoOpt is a collaborative and parallelized framework that comprises an open data platform and $K$ participants, each equipped with a distinct prior model. Specifically, CoOpt operates through the following five steps:
Step 1: Data distributing. The open data platform initiates the process by randomly partitioning the entire set of unlabeled data $D$ into $K$ non-overlapping subsets. Each participant then downloads one of these subsets from the platform, denoted as $D^{(k)}$, where $k$ denotes the $k$-th participant.
Step 2: Data optimization. Each participant optimizes their respective unlabeled dataset $D^{(k)}$ using the prior model $\boldsymbol{\psi}^{k}$. This data optimization process, defined in Proposition 1, yields optimized targets $D^{(k)}_{Y}$ and optimized data ${D^{\prime}}^{(k)}$.
Step 3: Data alignment. The heterogeneity of prior models among participants induces significant variations in their distribution of the targets. This issue, referred to as target distribution inconsistency (defined in Section 3.3), necessitates an alignment strategy (detailed in Section 3.4) to align the target distribution spaces across participants. Crucially, each participant needs to align their targets to the distribution space of the most optimal prior model using a learnable transformation matrix $\mathbf{T}^{(k)}$, yielding the final optimized dataset ${D^{\prime}}^{(k)}$. Further details are provided in Section 3.4.
Step 4: Data uploading. After optimization and alignment, participants upload their optimized datasets $\{{D^{\prime}}^{(k)}\}_{k=1}^{K}$ back to the open data platform.
Step 5: Data merging. The platform aggregates all the optimized datasets received from the participants to form a consolidated dataset.
The proposed collaborative parallel process enables participants to independently and parallelly optimize their subsets while ensuring consistency through the proposed alignment strategies. Consequently, this approach markedly reduces individual data optimization costs and enhances data processing efficiency through parallel execution.
3.3 An Inherent Challenge: Target Distribution Inconsistency
In our collaborative framework, each participant may employ a distinct prior model, leading to inconsistencies in the target distributions, as illustrated in Figure 5(a). For example, participant 1 uses ResNet-18 for optimization, resulting in a target dimension of $512$, while participant 2 utilizes ResNet-50, yielding a target dimension of $2{,}048$. Such inconsistencies can negatively impact the generalization capabilities of models trained on the optimized data, as they prevent the models from learning representations that are uniformly representative of the overall data distribution.
3.4 An Effective Strategy: Target Alignment
To address this issue, a potential solution is to align the target distributions of all participants’ prior models with that of the prior model producing the most optimal target distribution space, referred to as the best prior model. Such alignment can be achieved by utilizing an optimizable transformation matrix to map each participant’s target distribution to that of the best prior model [29]. This alignment strategy ensures consistency across all optimized target distribution spaces.
In summary, it is crucial to first effectively assess each participant’s prior model quality and subsequently train the transformation matrix for alignment.
A metric to quantify prior model quality.
Drawing inspirations from [32], which proposes an optimizable metric a.k.a. uniform value loss to achieve feature uniformity on the hypersphere during training, we employ this metric to evaluate the quality of prior models. Notably, [32] also provide theoretical validation of the connection between uniformity and feature quality. Specifically, each participant downloads a small shared dataset $S_{X}$ from the platform and computes the uniformity value of their prior model on $S_{X}$. They then upload this value to the platform, enabling it to determine which participant possesses the best prior model. The uniform value is computed as:
$\mathcal{V}_{\text{uniform}}(\boldsymbol{\psi};S)\triangleq\log\mathbb{E}_{\mathbf{x}_{i},\mathbf{x}_{j}\sim S}\left[e^{\tau\|{\boldsymbol{\psi}}(\mathbf{x}_{i})-{\boldsymbol{\psi}}(\mathbf{x}_{j})\|_{2}^{2}}\right],$ (3)
where $\boldsymbol{\psi}$ is the prior model, $\tau$ is a hyper-parameter set as $2$, consistent with [32].
A lower uniform value indicates a higher-quality prior model, which optimizes targets of superior quality. Extensive experiments in Figure 3(c) demonstrate a strong correlation between this metric and the performance of prior, thereby effectively assessing the quality of targets.
Alignment.
Upon identifying the best prior model, denoted as $\boldsymbol{\psi}^{\star}$, all participants, excluding the best prior model itself, proceed to train an optimizable transformation matrix. Specifically, the participant owning the best prior model computes its optimized targets, denoted as ${\mathbf{S}_{Y}}^{\star}$, onthe shared dataset $S_{X}$ using the prior model ${\boldsymbol{\psi}}^{\star}$. ${\mathbf{S}_{Y}}^{\star}$ are then uploaded to the platform, which ensures they are publicly accessible to all participants. Following this, each participant $k$ optimizes a lightweight transformation matrix, denoted as $\mathbf{T}^{(k)}$, on the shared dataset $S_{X}$. The optimization problem is defined as follows:
$\mathbf{T}^{(k)}=\operatorname*{arg\,min\,}_{\mathbf{T}\in\mathbb{R}^{n\times n}}\{\|\mathbf{T}\cdot{\boldsymbol{\psi}}^{(k)}(\mathbf{S}_{X})-{\mathbf{S}_{Y}}^{\star}\|_{2}^{2}\}\,,$ (4)
where $\mathbf{S}_{X}$ represents the matrix form of $S_{X}$, suitable for input into the network ${\boldsymbol{\psi}}^{(k)}$, and ${\mathbf{S}_{Y}}^{\star}$ also represents the matrix form of ${S_{Y}}^{\star}$. After obtaining the transformation matrix $\mathbf{T}^{(k)}$, the participant can convert the optimized targets for its own data using this matrix: ${D_{Y}}^{(k)}=\mathbf{T}^{(k)}\cdot{\boldsymbol{\psi}}^{(k)}({\mathbf{D}_{X}}^{(k)})$, where ${\mathbf{D}_{X}}^{(k)}$ denotes the participant’s subset, and ${D_{Y}}^{(k)}$ are the adjusted targets aligned with the best prior model’s target distribution space. As illustrated in Figure 5(b), the proposed alignment strategy effectively mitigates target distribution space inconsistency.
Remark on Privacy.
Notably,this work focuses on optimizing large-scale open-source unlabeled data obtained from publicly available sources, such as the internet. The information transmitted between the platform and participants is targets generated by prior models, which ensures that no direct privacy-sensitive information is exposed. Nevertheless, enhancing mechanisms for robust privacy protection remains a central objective for our future research.
Remark on Theoretical Effectiveness.
Our proposed framework builds upon well-established theoretical foundations. Specifically, RELA [29] has theoretically demonstrated that leveraging task-agnostic models, such as pre-trained models, can accelerate learning. Notably, our framework diverges substantially from RELA in both its objectives and methodology. Specifically, CoOpt introduces a collaborative optimization approach tailored for unlabeled data, targeting a fundamentally distinct problem domain. The challenges associated with collaborative data optimization are unique to CoOpt and remain unaddressed by RELA. Furthermore, the metric we employ to evaluate the quality of prior models, uniform value loss [32], has been theoretically validated for its effectiveness.
Table 2: Comparison of CoOpt with Self-Supervised Learning Methods in Accuracy (%) and Training Time (s). We use four datasets: CF-10 (CIFAR-10), CF-100 (CIFAR-100), T-IN (Tiny-ImageNet), and IN-1K (ImageNet-1K). The best results are marked in bold. $\uparrow$ means the performance improvement over the second-best result. $\times$ denotes the factor of training speed compared to the second-best result.
Dataset Metric BYOL DINO MoCo SimCLR SimSiam SwAV DCL CoOpt
CF-10 Acc. (%) 82.8 $\pm$ 0.1 82.6 $\pm$ 0.0 82.9 $\pm$ 0.1 83.1 $\pm$ 0.0 79.0 $\pm$ 0.0 82.9 $\pm$ 0.1 83.9 $\pm$ 0.1 88.7 $\pm$ 0.1 ($\uparrow$ 4.8)
Time (s) 1,376.56 1,457.22 1,349.56 1,114.81 1,090.79 1,012.74 1,783.34 540.43 ($\times$ 1.87)
CF-100 Acc. (%) 51.7 $\pm$ 0.1 51.0 $\pm$ 0.0 57.8 $\pm$ 0.1 55.4 $\pm$ 0.0 44.6 $\pm$ 0.1 53.2 $\pm$ 0.1 58.2 $\pm$ 0.2 65.8 $\pm$ 0.1 ($\uparrow$ 7.6)
Time (s) 1,406.17 1,419.69 1,425.80 1,103.45 1,139.14 1,072.44 1,701.49 548.11 ($\times$ 1.95)
T-IN Acc. (%) 43.9 $\pm$ 0.2 36.1 $\pm$ 0.0 42.4 $\pm$ 0.2 41.5 $\pm$ 0.1 40.8 $\pm$ 0.0 39.9 $\pm$ 0.1 44.6 $\pm$ 0.0 58.2 $\pm$ 0.1 ($\uparrow$ 13.6)
Time (s) 7,086.62 7,030.90 7,133.98 5,621.33 5,531.92 5,540.96 9,201.51 2,852.67 ($\times$ 1.94)
IN-1K Acc.(%) 61.9 $\pm$ 0.1 52.2 $\pm$ 0.0 57.6 $\pm$ 0.0 58.0 $\pm$ 0.0 55.8 $\pm$ 0.1 57.2 $\pm$ 0.1 60.6 $\pm$ 0.1 68.7 $\pm$ 0.0 ($\uparrow$ 6.8)
Time (s) 133,766.19 133,156.88 150,420.36 99,176.29 98,656.57 96,134.98 102,450.84 80,096.43 ($\times$ 1.20)
4 Experiments
In this section, we outline the experimental setup in Section 4.1 and assess the performance of our proposed collaborative data optimization framework, CoOpt, across various real-world scenarios, employing different datasets and architectures, as discussed in Section 4.2. Subsequently, we explore a continuous data optimization framework that allows the prior models of participants to evolve, thereby further optimizing the targets, as detailed in Section 4.3. Finally, we validate the effectiveness of the introduced uniform value metric and the target alignment strategy in Section 4.4.
4.1 Experimental Setup
Datasets and Networks:
We conduct experiments on both large-scale and small-scale datasets, including the full $224\times 224$ ImageNet-1k (224 $\times$ 224) [10], Tiny-ImageNet (64 $\times$ 64) [21], CIFAR-100 [19] and CIFAR-10 (32 $\times$ 32) [20]. Following previous self-supervised studies [15, 6, 13, 8, 1, 40], we employ a range of model capacities backbone architectures to evaluate the generalizability of our method, including ResNet-18, 50, 101 [16], ViT [12], and a series of CLIP-based models [26].
Baselines:
For the unlabeled data, referring to a prior widely-used benchmark [9], we consider several state-of-the-art self-supervised methods as baselines for a broader practical impact, including: SimCLR [6], BYOL [13], DINO [4], MoCo [15], SimSiam [8], SwAV [3], and DCL [36]. Notably, we do not compare our approach with knowledge distillation (KD) or dataset distillation (DD) methods, as summarized in Table 1, since they primarily foucus on the optimization of labeled data. In contrast, this work focuses on the utilization of unlabeled data.
Evaluation and Metrics:
Following previous benchmarks [15, 6, 13, 8], we evaluate the representation quality of the trained models by evaluating their test accuracy (%) using an offline linear probing strategy. Additionally, computational efficiency quantified by time cost (s).
Implementation details:
The proposed algorithm, CoOpt, involves an open data platform and facilitates interaction among multiple participants. The implementation follows five key steps (detailed in Section 3.2), and more details are provided in Appendix B.
  1. 1.
    For step 1: the training dataset is evenly distributed among all participants.
  2. 2.
    For steps 2 and 3, in practical applications, each participant can use publicly pre-trained models or their own models directly as the prior model. To simulate the diversity of prior models in practical applications, we utilize pre-trained models trained on ImageNet as well as a series of pre-trained CLIP-based models.
  3. 3.
    Steps 4 and 5 involve uploading data to the open platform for aggregation. Subsequently, for training on optimized data, we use the AdamW optimizer, the same as baselines. The size of mini-batch is set as $128$, except for ImageNet-1K, where a mini-batch size of 256 is utilized.
All experiments are conducted using 4 NVIDIA RTX 4090 GPUs. For all experiments, we utilize 3 random seeds and report both the mean and variance of the results. For fair comparisons, all methods in the experiments are executed with the same hyperparameters.
4.2 Comparison with the state-of-the-art methods
We evaluate our framework, CoOpt, across various scenarios: (1) participants leverage a diverse array of pre-trained prior models, characterized by variations in datasets, architectures, and training paradigms. (2) we specifically evaluate the robustness of our method when employing different prior datasets used for training prior models; (3) finally, we explore special cases involving human or weak models, particularly where there are resource-rich or resource-poor participants.
Diverse Pre-trained Prior Models.
Given the diversity of prior models, we employ a comprehensive set of pre-trained models, including those trained on ImageNet-1K and CLIP-based models. Details of these models are provided in Appendix B.2. These models are trained using various training paradigms, datasets, and architectures. As shown in Table 2, our method CoOpt demonstrates superior performance and training efficiency compared to existing self-supervised learning methods. Additional experimental results for the optimized data, trained across a wider range of architectures, are presented in Table 5 within Appendix B.3.
  1. (a)
    Performance: CoOpt significantly outperforms state-of-the-art self-supervised learning methods. For instance, it achieves an improvement of 13.6% and 6.8% over the leading self-supervised approach BYOL on Tiny-ImageNet and ImageNet-1K, respectively. Notably, CoOpt demonstrates more significant improvements on larger-scale datasets, such as Tiny-ImageNet and ImageNet-1K, which is particularly advantageous given the current emphasis on large-scale data era.
  2. (b)
    Efficiency: CoOpt demonstrates a substantial reduction in training costs across various datasets. Notably, on the Tiny-ImageNet, CoOpt achieves a training speed that surpasses BYOL by a factor of approximately ×$\times 1.94$.
Table 3: Comparison of CoOpt with BYOL Across Diverse Prior Datasets. For instance, “CIFAR-10 (P)” indicates participants’ prior models are trained on CIFAR-10. Bold means the best results. Underline indicates the results when the prior dataset is identical to the training data. All models are based on ResNet-18.
Dataset BYOL (Baseline) Our CoOpt (Diverse Prior Datasets)
CIFAR-10 (P) CIFAR-100 (P) Tiny-ImageNet (P) ImageNet-1K (P)
CIFAR-10 82.8 $\pm$ 0.1 86.6 $\pm$ 0.0 ($\uparrow 3.8$) 80.9 $\pm$ 0.0 ($\downarrow 1.9$) 81.6 $\pm$ 0.1 ($\downarrow 1.2$) 88.1 $\pm$ 0.0 ($\uparrow$ 5.3)
CIFAR-100 51.7 $\pm$ 0.1 54.9 $\pm$ 0.1 ($\uparrow 3.2$) 60.0 $\pm$ 0.1 ($\uparrow 8.3$) 56.8 $\pm$ 0.0 ($\uparrow 5.1$) 63.7 $\pm$ 0.0 ($\uparrow$ 12.0)
Tiny-ImageNet 43.9 $\pm$ 0.2 38.3 $\pm$ 0.0 ($\downarrow 5.6$) 40.2 $\pm$ 0.1 ($\downarrow 3.7$) 49.0 $\pm$ 0.0 ($\uparrow 5.1$) 55.8 $\pm$ 0.1 ($\uparrow$ 11.9)
ImageNet-1K 61.9 $\pm$ 0.1 31.7 $\pm$ 0.1 ($\downarrow 30.2$) 31.8 $\pm$ 0.0 ($\downarrow 30.1$) 40.5 $\pm$ 0.0 ($\downarrow 21.4$) 71.2 $\pm$ 0.0 ($\uparrow$ 9.3)
Diverse Prior Datasets.
In the last experiment, we employ pre-trained models, some of which are congruent with the unlabeled training data, while others are incongruent. We refer to these datasets as the prior datasets. To rigorously evaluate the influence of these prior datasets, we perform an analysis across scenarios where the prior datasets used for prior models either align with or differ from the unlabeled training dataset. For instance, in the aligned scenario, the training dataset is CIFAR-10, and the prior models are also trained on CIFAR-10 (CIFAR-10 (P)). Conversely, in the divergent scenario, the training dataset remains CIFAR-10, while the prior models are trained on CIFAR-100 (P). Details of these models are provided in Appendix B.4. We evaluate our approach on four publicly available datasets, with the results summarized in Table 3.
  1. (a)
    Utilizing prior models trained on the same dataset as the target training data can lead to notable performance gains.
  2. (b)
    For complex training datasets, leveraging prior models trained on simpler datasets may degrade performance compared to BYOL This degradation arises due to the limited informativeness of simpler prior datasets, which provide weaker guidance.
  3. (c)
    Importantly, for all unlabeled training datasets, employing prior models trained on ImageNet-1K consistently yields significant performance gains. This can be attributed to the strong generalization abilities of such models. This observation is particularly pertinent in practical applications, as the majority of publicly available pre-trained models are derived from ImageNet-1K.
Table 4: Comparison of CoOpt with BYOL in Presence of Human or Weak Prior Models.
Prior Models Dataset
Method Human Weak CIFAR-10
BYOL 82.8 $\pm$ 0.1
CoOpt 88.7 $\pm$ 0.1
89.0 $\pm$ 0.0
90.3 $\pm$ 0.0
89.2 $\pm$ 0.0
Extreme Cases of Prior Models: Human or Weak.
In real-world applications, extreme cases arise due to the varying capabilities of participants. For example, some participants have extensive resources and employ human annotators for labeling, while others may have limited resources and rely on weak models with inferior performance and generalization abilities. In our experiments, we define weak models as those trained during intermediate stages that are even far from convergence. To simulate the conditions, in addition to the prior models used in the first experiment, we incorporate $5$ prior models, either human or weak models to the data optimization process. The results are summarized in Table 4. Surprisingly, even the inclusion of weaker models contributes to enhancing the final performance, indicating that such models can still provide valuable information. Moreover, it is important to note that the integration of high-capacity, human-like models results in significant performance improvements.
(a) Training curves
(b) Varying Prior Models
(c) Uniform value
(d) Shared data size
Figure 3: Comprehensive Analysis of CoOpt. (a) Training curves: Comparison of various self-supervised learning methods and the proposed CoOpt. (b) Prior Models With Varying Accuracies. Employing a strong model as the prior model is not essential (c) Correlation Verification: Examine the relationship between the uniform value and performance, which demonstrates a strong negative correlation. (d) Influence of shared data size: As the size of the shared data increases, the performance gains diminish.
Moreover, we conduct experiments with only weaker model, as shown in Figure 3(b). Here, The dashed line represents the training curve of BYOL (baseline), while the solid lines correspond to prior models with different accuracies. "Prior model (BYOL)" indicates the use of BYOL as the prior model, and the accuracies of the other prior models are all lower than that of BYOL. Notably, most prior models surpass BYOL performance in the early stages of training, and even a model with approximately 75% accuracy can outperform BYOL in the final performance. The results demonstrate that employing a strong model as the prior model is not essential.
Figure 4: A Practical Scenario: Continuous Optimization Prior models experience temporal evolution, leading to improved target quality across multiple rounds.
4.3 Continuous Data Optimization
We explore another practical scenario where the prior models undergo temporal evolution. For example, a participant’s initial model, such as ResNet-50, might be upgraded to a higher-capacity model like ResNet-101 as their computational resources improve. Consequently, the process can be treated as a dynamic and continuous procedure. The detailed description of the process is provided in Appendix B.5.
We simulate this scenario by conducting experiments where, in each interaction round between the platform and participants, $20$% of the participants randomly update their model architectures to reflect an increase in model capacity. The training curves across 10 rounds on CIFAR-100 are depicted in Figure 4. The results have demonstrated that in CoOpt, as the prior models evolve, the quality of the targets improves, thereby facilitating continuous optimization. In particular, compared to the baseline BYOL, these improvements can achieve an enhancement of $4.6$%.
4.4 Ablation Study
Effectiveness of Uniform Value.
To evaluate the effectiveness of uniform value in estimating the quality of prior models, we employ a diverse set of prior models, calculating both their uniform value and test accuracy. We employ the Spearman rank correlation coefficient [^4] The formula is: $\rho=1-\frac{{6\sum d_{i}^{2}}}{{n(n^{2}-1)}}$, where $d_{i}$ represents the difference between the ranks of each pair of observations and $n$ denotes the number of observations. $\rho$ [39] to quantify the association between Uniform Value and accuracy. As illustrated in Figure 3(c), the Spearman rank correlation coefficient is $\rho=-0.9714$, indicating a strong correlation between uniform value and model performance, thereby effectively assessing the quality of prior models.
Target Distribution Alignment.
In the real-world applications, participants often employ diverse prior models, which results in target distribution space inconsistencies, as demonstrated in Figure 5(a). We conduct an ablation study w/o alignment and w/ alignment to demonstrate the importance and necessarily of target distribution space alignment, and the training curves presented in Figure 5(d). Three key observations include:
  1. (a)
    Compared to without alignment (green line), our proposed approach (blue line) yields a performance improvement of $16.9$%, which highlights the critical importance of aligning the target distribution space. Notably, the training curve without alignment (green line) initially ascends before declining, suggesting that during the early phase of model training, optimizing the target leads to performance enhancement. However, as training continues, the severely inconsistent targets significantly degrade model performance.
  2. (b)
    Furthermore, we compare our alignment strategy, which aligns to the best prior model, with two other straightforward strategies: aligning to a medium prior model (purple line) and aligning to the worse prior model (red line). Obviously, three strategies outperform the scenario without alignment, and aligning to the best prior model provides the most significant performance gains.
  3. (c)
    To further analyze the underlying reasons, we employ t-SNE visualization. Notably, a comparison between alignment with the better participant ( Figure 5(b)) and the worse participant ( Figure 5(c)) reveals that alignment with a worse-quality model diminishes the representative capability of the targets. This, in turn, results in less effective guidance for model training.
Influence of Shared Unlabeled Data Size.
The shared unlabeled data $S$ is used to estimate the uniform value and compute the transformation matrix for target alignment, as detailed in Section 3.4. To explore the influence of the size of shared unlabeled data, we conduct experiments on both CIFAR-10 and CIFAR-100, varying the size of the shared data in proportions of $\{0.01,0.05,0.1,0.2,0.4,0.6,0.8\}$ relative to the original dataset. The results are presented in Figure 3(d). It is observed that as the size of the shared data increases, performance gains become marginal. Specifically, increasing the proportion of data from $0.01$% to $0.05$% results in improvements of $2.4$% and $2.5$% for CIFAR-10 and CIFAR-100, respectively. However, increasing the proportion from $0.1\%$ to $0.8\%$ results in only a $0.5\%$ and $0.6\%$ improvement. This suggests that a small-sized shared unlabeled dataset is adequate for accurate uniform value estimation and the computation of the transformation matrix.
(a) Before Alignment
(b) Align (BYOL $\rightarrow$ CE)
(c) Align (CE $\rightarrow$ BYOL)
(d) Training curves.
Figure 5: Effectiveness of Target Distribution Alignment. (a), (b), (c): Visualization of t-SNE for optimized targets generated by two distinct models (BYOL (acc. = 82%) and SupCE (acc. = 90%).) Aligning to the worse model (c) results in diminished target quality. (d): Training curves with and without alignment.
5 Conclusion and Future Work
Conclusion.
In this paper, we introduce CoOpt, a pioneering data-centric, efficient, and parallelized framework for collaborative optimization of unlabeled data. This data-centric approach results in architecture-agnostic optimized data that are reusable across diverse network architectures, while simultaneously reducing the number of training iterations required, thereby enhancing overall computational efficiency. The collaborative mechanism effectively exploits distributed computation to substantially lower the cost per participant. Furthemore, within CoOpt, we identify a critical issue: Target Distribution Inconsistency, which arises from the diversity of prior models used in data optimization. To mitigate this, we propose a target alignment strategy. Extensive experiments conducted across various real-world scenarios demonstrate the superior effectiveness and efficiency of the CoOpt framework across diverse datasets and architectures. One limitation is that the experiments are primarily conducted on image datasets, without extending to other domain data.
Future Work.
In the future, our objectives include the following: (1) Develop advanced methodologies to ensure strong privacy guarantees; (2) Conduct an in-depth exploration of continuous data optimization through iterative processes, enabling dynamic collaboration among participants. (3) Integrate diverse domain datasets into the optimization framework. (3) Explore more practical scenarios, such as participants optimizing their local data and then uploading it to the shared platform. This paradigm can broaden the platform’s functionalities, where organizations and individuals can contribute to and benefit from a globally data ecosystem.
Appendix A Related Work
Self-supervised Learning: A Model-Centric Perspective.
Self-supervised learning [7] aims to exploit the intrinsic relationships within unlabeled data. For example, InstDisc [34] uses instance discrimination as a pretext task. Building on this, CMC [30] proposes to use multiple views of an image as positive samples and take another one as the negative. MoCo [15] significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples. SimCLR [6] highlights the importance of hard positive sample strategies by introducing data augmentation. Notably, BYOL [13] discards negative sampling and surpasses the performance of SimCLR [6].
Summary. They are constrained to specific architectures and incur high computational costs due to the reliance on large batch sizes or memory banks.
Knowledge Distillation: Optimzing Targets.
Knowledge distillation [17] employs soft labels generated by teacher models to improve the performance of a student model and expedite its training [11]. Many following works aim to enhance the use of soft labels for more effective knowledge transfer. For example, WSLD [45] analyzes soft labels and distributes different weights for them from a perspective of bias-variance trade-off. DKD [44] decouples the logits and assigns different weights for the target and non-target classes. Moreover, several studies [37, 11] demonstrate that knowledge distillation can accelerate the training process.
Dataset Distillation: Optimizing Both Samples and Targets.
Dataset distillation [33] aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost. The majority of methods focus on optimizing images, which can be categorized into five primary approaches [22]: meta-learning frameworks [33, 46], gradient matching [43, 41], distribution matching [42, 38], trajectory matching [5, 14], and decoupling frameworks [38, 29]. Recently, some studies [27, 25] have shifted focus towards label distillation, aiming to obtain high-quality soft labels. This approach has demonstrated notable efficiency and effectiveness.
Summary. Both approaches are computationally expensive as they require task-specific models.
Appendix B Experimental Details
B.1 Implementation Details
Hardware Setup.
All experiments are conducted using 4 NVIDIA RTX 4090 GPUs. For fair comparisons, all methods in the experiments are executed with the same hyperparameters.
Unlabeled Training Dataset Split.
In our framework, for all experiments, the unlabeled dataset is evenly distributed among all participants.
B.2 Diverse Pre-trained Prior Models.
We utilize 10 pre-trained prior models, which are downloaded using the torchvision package. These models consist of 5 ImageNet-1K-based models, varying from small-scale to large-scale architectures, and 5 CLIP-based model. Unlabeled data is equally split among participants.
B.3 Diverse Network for Model training in Optimized data.
Table 5: Comparison of CoOpt with BYOL for Diverse Network on CIFAR-100.
Method
Network BYOL CoOpt
ResNet-50 60.4 $\pm$ 0.1 63.8 $\pm$ 0.0
ResNet-101 61.5 $\pm$ 0.2 65.7 $\pm$ 0.2
MobileNet-v2 24.0 $\pm$ 0.5 58.1 $\pm$ 0.0
Efficientnet-b0 2.3 $\pm$ 0.2 70.7 $\pm$ 0.2
ViT 38.5 $\pm$ 0.1 57.8 $\pm$ 0.1
We conduct experiments across various network architectures, comparing our approach, CoOpt, with BYOL, as presented in Table 5. Notably, CoOpt consistently demonstrates significant performance improvements over BYOL across multiple architectures. Furthermore, BYOL exhibits markedly inferior performance when applied to lightweight architectures such as MobileNet-v2 and EfficientNet-b0, in contrast to its performance with larger backbones like ResNet-50. One plausible reason for this performance drop is the instability of batch normalization (BN) statistics in the early layers of these lightweight networks. Specifically, the narrow layers (i.e., those with small channel width) in early stages produce activations with limited sample size per channel, resulting in high variance in the estimated batch mean and variance. Since BYOL relies on two differently augmented views of the same image, this BN-induced instability leads to exaggerated discrepancies between the two projected features, which deviates from the intended invariance objective of the method. This mismatch potentially impairs the optimization process and harms representation quality.
B.4 Diverse Prior Datasets
In this set of experiments, we investigate the effectiveness of our proposed CoOpt when the prior datasets are either similar to or distinct from the training datasets. For each prior dataset, we train 4 different models to serve as prior models. Specifically: These 4 models are trained using two paradigms (supervised and unsupervised learning) and two architectures (ResNet-18 and ViT). For each training dataset, the data is evenly distributed among participants. Each prior model is assigned 1/4 of the unlabeled dataset, which it optimizes independently. The results are then aggregated on the open platform.
B.5 Continuous Data Optimization
Figure 6: Common Target Conflict Scenarios
In real-world scenarios, model architectures and computational resources often evolve over time, necessitating a dynamic and continuous approach to data optimization. Unlike static optimization, where data is processed only once, this framework accommodates temporal model evolution and repeatedly refines optimized data to achieve superior results. Below, we describe the core components of this framework.
Dynamic Model Evolution.
Initially, all data is optimized using participants’ current models. Over time, as computational resources improve, participants upgrade to higher-capacity architectures. These upgraded models exhibit stronger feature extraction capabilities, enabling further refinement of previously optimized data. This dynamic evolution transforms the optimization process into a continuous improvement cycle.
Open Platform for Optimized Data Comparison.
A key feature of this framework is the inclusion of an open platform for optimized data submission and evaluation. Participants provide newly optimized data, which is compared against previously optimized data to retain the superior results. This comparison leverages evaluation metrics such as the uniform value of the prior model, ensuring that the dataset evolves toward higher optimization quality over successive iterations.
Iterative Optimization Process
The overall process is illustrated in Figure 6. The framework operates as a loop:
  1. 1.
    Data is initially optimized by participants’ models.
  2. 2.
    As models evolve, data is re-optimized to reflect the improved capacities of the upgraded architectures.
  3. 3.
    The open platform compares new and old optimization results, retaining the higher-quality data.
  4. 4.
    This process repeats over multiple interaction rounds, progressively enhancing the dataset.

### v2

@@
-Collaborative Unlabeled Data Optimization
+Beyond Model-centric: Collaborative DATA Optimization for Reusing and Sharing
 Abstract
-This paper pioneers a novel data-centric paradigm to maximize the utility of unlabeled data, tackling a critical question: How can we enhance the efficiency and sustainability of deep learning training by optimizing the data itself?
-We begin by identifying three key limitations in existing model-centric approaches, all rooted in a shared bottleneck: knowledge extracted from data is locked to model parameters, hindering its reusability and scalability.
-To this end, we propose CoOpt, a highly efficient, parallelized framework for collaborative unlabeled data optimization, thereby effectively encoding knowledge into the data itself.
-By distributing unlabeled data and leveraging publicly available task-agnostic models, CoOpt facilitates scalable, reusable, and sustainable training pipelines.
-Extensive experiments across diverse datasets and architectures demonstrate its efficacy and efficiency, achieving 13.6% and 6.8% improvements on Tiny-ImageNet and ImageNet-1K, respectively, with training speedups of 1.94×1.94$\times$1.94 × and 1.2×1.2$\times$1.2 ×  [^1] Our code is provided in the Supplementary Materials and will be publicly accessible..
+This paper pioneers a novel data-centric paradigm to maximize the utility of unlabeled data, tackling a critical question: How can we enhance the sustainability and efficiency of deep learning training by optimizing the data itself?
+We begin by identifying two key limitations in existing model-centric approaches, all rooted in a shared bottleneck: knowledge extracted from data is locked to model parameters, hindering its reusability and scalability.
+To this end, we propose CoOpt, a highly efficient, parallelized framework for collaborative unlabeled data optimization.
+By distributing unlabeled data and leveraging publicly available task-agnostic prior models, CoOpt optimizes raw unlabeled data into knowledge-enriched training sets that are effective, efficient, reusable, and easily shareable.
+Extensive experiments across diverse datasets and architectures validate these advantages, achieving a 7.9% improvement on ImageNet-1K over BYOL.
+Notably, CoOpt remains effective even when all prior models are significantly weak, substantially accelerating the early stages of training.
+These results establish data-centric optimization as a promising path toward sustainable and efficient deep learning  [^1] Code will be publicly available..
 1 Introduction
@@
 Therefore, we propose CoOpt, an Refer to caption efficient and parallel framework enabling participants to use diverse task-agnostic models, such as pre-trained ResNets, termed prior models, for collaborative data optimization.
-Deep Learning has achieved remarkable success, primarily due to the large-scale datasets [28, 35].
-Despite the abundance of data in the era of big data, a significant portion of them remains unlabeled [22].
-The dominant paradigm in the field for exploiting unlabeled data is self-supervised learning (SSL), which is fundamentally model-centric: it carefully crafted pretext tasks and loss functions to encode data information into model parameters [6, 13, 4].
-However, the model-centric nature leads to three fundamental challenges.
-First, their training protocols are tightly coupled to specific network architectures, significantly hindering the transferability of trained model to other architectures [31, 18].
-Second, even with advances in acceleration techniques, training over extensive unlabeled datasets still incurs prohibitively high computational costs [29].
-Third, when a new downstream task arises, pretrained models typically require extensive fine-tuning or even full retraining.
+Deep Learning has achieved remarkable success, primarily due to the large-scale datasets (Song et al., 2020; Yang et al., 2023).
+Despite the abundance of data in the era of big data, a significant portion of them remains unlabeled (Lei & Tao, 2023).
+The dominant paradigm in the field for exploiting unlabeled data is self-supervised learning (SSL), which is fundamentally model-centric: it carefully crafted pretext tasks and loss functions to encode data information into model parameters (Chen et al., 2020a; Grill et al., 2020; Gui et al., 2024).
+However, the model-centric nature presents two critical challenges.
+First, their training protocols are tightly coupled to specific network architectures, severely hindering the transferability and reusability of trained model to other architectures (Wagner et al., 2022; Huang et al., 2023).
+Second, despite acceleration advances, training over extensive unlabeled datasets still computationally prohibitive (Sun et al., 2024).
 At the core of these challenges is a shared bottleneck: knowledge extracted from data is locked in model parameters, restricting its adaptability and preventing efficient reuse across diverse tasks or architectures.
-To break free from model-centric paradigm, we propose a data-centric paradigm that directly optimizes the unlabeled data by optimizing targets for samples (further elaborated in Section 3.1), thereby effectively encoding knowledge into the data itself rather than into model parameters.
+To break free from model-centric paradigm, we propose a data-centric paradigm that directly optimizes the unlabeled data by optimizing targets for samples (detailed in Sec.
+3.1), thereby effectively encoding knowledge into the data itself rather than into model parameters.
 The resulting “optimal data” is agnostic to downstream architectures, accelerates subsequent training by providing richer supervision, and can be reused across multiple tasks without repeated large-scale pretraining.
@@
 To address this, we propose CoOpt, a highly efficient collaborative framework inspired by crowd-sourcing to achieve parallel data optimization.
-An overview of CoOpt is depicted in Figure 1, with detailed processes shown in Figure 2.
+An overview of CoOpt is depicted in Fig.
+1, with detailed processes shown in Fig.
+3.
 In CoOpt, the unlabeled dataset is partitioned into disjoint subsets, each processed independently and in parallel by participants equipped with task-agnostic models.
@@
 Once the targets are optimized, they are aggregated to reconstruct a fully optimized dataset, achieving computational scalability through decentralized workload distribution.
-CoOpt offers three key advantages: First, as the optimization relies solely on task-agnostic models, the aggregated optimized targets can be directly applied to any downstream architecture.
-Second, by distributing data subsets across participants, each node incurs only a fraction of the total computational load, enabling scalability and resource efficiency.
-Importantly, since only targets $D_{Y}$ are optimized, the aggregated targets are the same as those obtained by optimizing the whole dataset.
-Third, downstream tasks can directly fine-tune the optimized dataset for task-specific objectives without requiring training from scratch.
-Comprehensive experimental results have demonstrated that CoOpt achieves superior performance, generalization, and training efficiency.
+CoOpt offers three key advantages: First, relying solely on task-agnostic prior models, the optimized data can be directly used to any downstream architecture, thereby ensuring strong generalizability and reusability (see Sec.
+4.2).
+Second, by distributing non-overlapping data subsets across participants, each node handles only a fraction of the total computational cost, thereby enabling scalable and resource-efficient optimization (see Sec.
+4.2).
+Third, CoOpt is lightweight, incurring negligible overhead while substantially enhancing training efficiency and performance (see Sec.
+4.4).
 In summary, our contributions are threefold:
   1. (a)
-    We propose CoOpt, the first data-centric framework for optimizing unlabeled data, leveraging distributed computation to significantly lower the computational and storage costs per participant.
+    We propose CoOpt, the first data-centric framework for collaboratively optimizing unlabeled data.
+By leveraging task-agnostic prior models, CoOpt transforms raw unlabeled samples into optimal data, enabling high performance, efficiency, strong generalization, and reusability.
   2. (b)
-    Within CoOpt, we identify a critical issue: Target Distribution Inconsistency in Section 3.3. Accordingly, we propose an effective target alignment strategy, elaborated in Section 3.4.
+    Within CoOpt, we identify a critical issue, Target Distribution Inconsistency (Sec. 3.3), and introduce a lightweight target alignment strategy to address it (Sec. 3.4).
   3. (c)
-    Extensive experiments have been conducted across a range of real-world scenarios, involving a variety of prior models trained on various datasets, architectures, and training paradigms.
-Notably, special cases are explored where human or significantly weak models are employed as prior models, verifying the robustness and flexibility of CoOpt.
-These experiments consistently demonstrate that CoOpt achieves superior effectiveness and efficiency.
+    We conduct experiments across datasets and models to comprehensively validate the advantages of CoOpt (Sec.
+4.2).
+Further, we provide a detailed analysis of the key factors influencing its effectiveness (Sec.
+4.3).
+Remarkably, we demonstrate that CoOpt remains effective even when all prior models are weak, substantially accelerating the early stages of training.
 2 Related Work
-Model-centric Perspective: Self-supervised Learning.
+Self-Supervised Learning.
 It aims to exploit the intrinsic relationships within unlabeled data.
-For example, InstDisc [34] uses instance discrimination as a pretext task.
-MoCo [15] significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples.
-SimCLR [6] highlights the importance of hard positive sample strategies.
-Notably, BYOL [13] discards negative sampling and surpasses the performance of SimCLR [6].
-Data-centric Perspective: Knowledge Distillation and Dataset Distillation.
-Knowledge distillation [17] employs soft labels generated by teacher models to improve the performance of a student model and expedite its training [11].
-Many following works [45, 44, 11] aim to enhance the use of soft labels.
-Dataset distillation [33] aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost.
-The majority of methods focus on optimizing images, which can be categorized into three primary approaches [22]: meta-learning frameworks [33, 46], matching-based methods [43, 42, 14] and decoupling frameworks [38, 29].
-A more detailed description of knowledge distillation and dataset distillation is provided in Section A.
+For example, InstDisc (Wu et al., 2018) uses instance discrimination as a pretext task.
+MoCo (He et al., 2020) significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples.
+SimCLR (Chen et al., 2020a) highlights the importance of hard positive sample strategies.
+Notably, BYOL (Grill et al., 2020) discards negative sampling and surpasses the performance of SimCLR (Chen et al., 2020a).
+Model-Centric Perspective: Knowledge Distillation.
+Knowledge distillation (Hinton, 2015) leverages teacher-generated soft labels to improve student training efficiency and performance (Dong et al., 2023).
+A line of knowledge distillation methods utilizes multiple teachers (MKD) (Zhang et al., 2022; Pham et al., 2023) to enhance student learning.
+They assume that ensemble outputs from multiple teachers enables students to learn more generalized representations.
+Notably, all teacher models process the same input data.
+Our setting departs fundamentally from knowledge distillation in terms of objective, input data, and teacher models (see Fig.
+2).
+First, in KD, the distilled knowledge is embedded in student parameters, limiting its reuse across different architectures.
+In contrast, our objective is to construct a high-quality, optimized dataset that is model-agnostic and reusable, enabling training or evaluation of diverse architectures.
+Second, rather than feeding all teachers the same inputs, we partition the unlabeled data into disjoint subsets, each optimized by a different prior model.
+Third, existing KD methods often rely on intricate loss functions (Jiang et al., 2024) or require teacher fine-tuning (Wu et al., 2021), but our framework uses arbitrary pre-trained models without domain-specific adaptation.
+The optimized data can then be directly reused to train arbitrary downstream models without further modification.
+Data-Centric Perspective: Dataset Distillation.
+Dataset distillation (Wang et al., 2018) improves the training efficiency by learning a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost.
+The majority of methods focus on optimizing images, which can be categorized into three primary approaches (Lei & Tao, 2023): meta-learning frameworks (Wang et al., 2018; Zhou et al., 2022), matching-based methods (Zhao et al., 2020; Zhao & Bilen, 2023; Guo et al., 2024) and decoupling frameworks (Yin et al., 2023; Sun et al., 2024).
+Notably, current methods predominantly focus on distilling labeled datasets.
 3 Collaborative Data Optimization Framework CoOpt
-We begin by formally defining data optimization in Section 3.1.
-Subsequently, we provide a detailed description of the proposed CoOpt in Section 3.2.
-Furthermore, we identify an inherent challenge within this framework in Section 3.3 and present method in Section 3.4.
-Table 1: Properties of Various Data Utilization Methods.
-‘Optimize $D_{X}$’ and ‘Optimize $D_{Y}$’ indicate whether the optimization process encompasses the samples and targets, respectively.
-‘SSL’ is self-supervised learning, ‘KD’ is knowledge distillation, and ‘DD’ is dataset distillation.
-‘$\mathbf{c}$’ denotes the training cost of standard supervised learning that employs human-labeled data.
-Method Optimize Efficiency Cost Analysis Total Efficiency
-$D_{X}$ $D_{Y}$
-Data
-Optimization
-Model
-Training
-Original SSL 0 [Uncaptioned image] ($\geq 2\mathbf{c}$) Extensive Computation [Uncaptioned image]
-Labeled Data Optimization KD [Uncaptioned image] ($\geq\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Task-specific Teacher Models [Uncaptioned image]
-DD [Uncaptioned image] ($\geq\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Task-specific Pre-trained models [Uncaptioned image]
-CoOpt (Our) [Uncaptioned image] ($<\mathbf{c}$) [Uncaptioned image] ($<\mathbf{c}$) Various Prior Models [Uncaptioned image]
+We begin by formally defining data optimization in Sec.
+3.1.
+Subsequently, we provide a detailed description of the proposed CoOpt in Sec.
+3.2.
+Furthermore, we identify an inherent challenge within this framework in Sec.
+3.3 and present method in Sec.
+3.4.
+Figure 2: Comparison Between KD and Ours.
 3.1 Definition of Data Optimization
-We first revisit current training acceleration techniques from a data-centric perspective, including knowledge distillation [17] and dataset distillation [33].
-Specifically, we decouple data into samples $D_{X}$ and targets $D_{Y}$  [^2] A formal definition of $D_{X}$ and $D_{Y}$ is provided in Section 3.1., as summarized in Table 1.
-Essentially, compared to the self-supervised learning methods, these approaches achieve higher efficiency by either optimizing targets $D_{Y}$ pre-trained models (as in knowledge distillation), or by optimizing both input data $D_{X}$ and targets $D_{Y}$ (as in dataset distillation).
-Notably, for dataset distillation, optimizing input data $D_{X}$ is computationally more expensive than optimizing targets $D_{Y}$ [2].
-Furthermore, recent studies [27, 25] have indicated that solely optimizing $D_{Y}$ not only reduces computational overhead but also achieves significant performance gains.
+We first revisit current training acceleration techniques, including knowledge distillation (Hinton, 2015) and dataset distillation (Wang et al., 2018).
+Specifically, we decouple data into samples $D_{X}$ and targets $D_{Y}$.
+Essentially, compared to self-supervised learning methods, these approaches are more efficient.
+They achieve this by either optimizing the target $D_{Y}$ of pre-trained models (as in knowledge distillation) or by jointly optimizing both the input data $D_{X}$ and target $D_{Y}$ (as in dataset distillation).
+Notably, for dataset distillation, optimizing input data $D_{X}$ is computationally more expensive than optimizing targets $D_{Y}$ (Bohdal et al., 2020).
+Furthermore, recent studies Shang et al.
+(2025); Qin et al.
+(2024) have indicated that solely optimizing $D_{Y}$ not only reduces computational overhead but also achieves significant performance gains.
 These findings have demonstrated that optimizing $D_{Y}$ is both necessary and efficient.
@@
 where $P_{\mathcal{T}}$ denotes the test distribution, $\mathbf{x}$ is a test sample, $y$ is its label, $\ell$ is the loss function (e.g., cross-entropy loss), and $\boldsymbol{\theta}_{D}$ and $\boldsymbol{\theta}_{D^{\prime}}$ are parameters of network $\boldsymbol{\phi}$ trained on $D$ and $D^{\prime}$, respectively.
-Notably, directly extending these methods to unlabeled data presents two challenges.
-First, they rely on task-specific models trained on labeled data, which are unavailable.
-Second, they incur substantial costs.
-For example, most dataset distillation methods rely on bi-level optimization [23].
+Notably, directly extending these methods to unlabeled data is infeasible.
+Existing dataset distillation methods focus on labeled data, whereas knowledge distillation methods, as reviewed in Sec.
+2, significantly differ from ours.
+A detailed comparison is presented in Fig.
+2.
 To address these limitations, we propose a collaborative framework that leverages distributed computation and various task-agnostic models for unlabeled data.
-Specifically, inspired by [29], which shows that using task-agnostic models for target assignment can expedite training, we further enhance efficiency by splitting the data and then parallelly optimizing each split.
+Specifically, inspired by (Sun et al., 2024), which shows that using task-agnostic models for target assignment can expedite training, we further enhance efficiency by splitting the data and then parallelly optimizing each split.
 After optimization, the optimal subset are aggregated to reconstruct a fully optimized dataset, achieving computational efficiency.
-Formally, we define data optimization with a prior model $\boldsymbol{\psi}$ in each participant in Proposition 1.
-{propositionframe}
-Proposition 1 (Data optimization with prior model $\boldsymbol{\psi}$) .
+Formally, we define data optimization with a prior model $\boldsymbol{\psi}$ in each participant in Def.
+1.
+{definitionframe}
+Definition 1 (Data optimization with prior model $\boldsymbol{\psi}$) .
 Given samples $D_{X}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ and a prior model $\boldsymbol{\psi}:\mathbb{R}^{m}\to\mathbb{R}^{l}$, data optimization assigns optimal targets $D_{Y}=\{\mathbf{y}_{i}\}_{i=1}^{N}$ for the samples to create $D^{\prime}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$.
@@
 where $\mathbf{y}_{i}$ is the optimized target, and $\boldsymbol{\psi}(\mathbf{x}_{i})$ represents the target of $\mathbf{x}_{i}$, which means the feature representation.
-$\mathbf{W}:\mathbb{R}^{l}\to\mathbb{R}^{n}$ denotes a matrix designed to transform the feature vector $\boldsymbol{\psi}(\mathbf{x}_{i})$ from dimension $l$ to $n$ without loss of information [24].
+$\mathbf{W}:\mathbb{R}^{l}\to\mathbb{R}^{n}$ denotes a matrix designed to transform the feature vector $\boldsymbol{\psi}(\mathbf{x}_{i})$ from dimension $l$ to $n$ without loss of information (Matoušek, 2008).
 This transformation aligns the output dimension [^2] Here, $n$ is the target dimensionality of $\boldsymbol{\phi}_{\boldsymbol{\theta}_{D^{\prime}}}$.
@@
 3.2 Overview of the Proposed Framework CoOpt
-CoOpt is a collaborative and parallelized framework that comprises an open data platform and $K$ participants, each equipped with a distinct prior model.
-Specifically, CoOpt operates through the following five steps:
+CoOpt is a collaborative and parallelized framework that includes an open data platform and $K$ participants, each equipped with a distinct prior model. CoOpt operates through the five steps:
 Step 1: Data Distributing.
@@
 Each participant optimizes their respective unlabeled dataset $D^{(k)}$ using the prior model $\boldsymbol{\psi}^{k}$.
-This data optimization process, defined in Proposition 1, yields optimized targets $D^{(k)}_{Y}$ and optimized data ${D^{\prime}}^{(k)}$.
+This data optimization process, defined in Prop.
+1, yields optimized targets $D^{(k)}_{Y}$ and optimized data ${D^{\prime}}^{(k)}$.
 Step 3: Data Alignment.
 The heterogeneity of prior models among participants induces significant variations in their distribution of the targets.
-This issue, referred to as target distribution inconsistency (defined in Section 3.3), necessitates an alignment strategy (detailed in Section 3.4) to align the target distribution spaces across participants.
-Crucially, each participant needs to align their targets to the distribution space of the most optimal prior model using a learnable transformation matrix $\mathbf{T}^{(k)}$, yielding the final optimized dataset ${D^{\prime}}^{(k)}$.
-Further details are provided in Section 3.4.
+This issue, referred to as target distribution inconsistency (defined in Sec.
+3.3), necessitates an alignment strategy (detailed in Sec.
+3.4) to align the target distribution across participants.
+Crucially, each participant needs to align their targets distribution to the most optimal prior model using a learnable transformation matrix $\mathbf{T}^{(k)}$, yielding the final optimized dataset ${D^{\prime}}^{(k)}$.
+Further details are provided in Sec.
+3.4.
 Step 4: Data Uploading. After optimization and alignment, participants upload their optimized datasets $\{{D^{\prime}}^{(k)}\}_{k=1}^{K}$ back to the open data platform.
 Step 5: Data Merging. The platform aggregates all the optimized datasets received from the participants to form a consolidated dataset.
-The proposed collaborative parallel process enables participants to independently and parallelly optimize their subsets while ensuring consistency through the proposed alignment strategies.
-Consequently, this approach markedly reduces individual data optimization costs and enhances data processing efficiency through parallel execution.
+The proposed CoOpt enables participants to independently and parallelly optimize their subsets while ensuring consistency through the proposed alignment strategies.
+Consequently, this approach markedly reduces individual data optimization costs and enhances efficiency.
 3.3 An Inherent Challenge: Target Distribution Inconsistency
-In our collaborative framework, each participant may employ a distinct prior model, leading to inconsistencies in the target distributions, as illustrated in Figure 5(a).
-For example, participant 1 uses ResNet-18 for optimization, resulting in a target dimension of $512$, while participant 2 utilizes ResNet-50, yielding a target dimension of $2{,}048$.
+In our collaborative framework, each participant may employ a distinct prior model, leading to inconsistencies in the target distributions, as illustrated in Fig.
+6(a).
+For example, participant 1 uses ResNet-18 for optimization, resulting in a target dimension of $512$, while participant 2 utilizes ResNet-50, yielding a target dimension of 2,0482{,}048.
 Such inconsistencies can negatively impact the generalization capabilities of models trained on the optimized data, as they prevent the models from learning representations that are uniformly representative of the overall data distribution.
 3.4 An Effective Strategy: Target Alignment
-To address this issue, a potential solution is to align the target distributions of all participants’ prior models with that of the prior model producing the most optimal target distribution space, referred to as the best prior model.
-Such alignment can be achieved by utilizing an optimizable transformation matrix to map each participant’s target distribution to that of the best prior model [29].
-This alignment strategy ensures consistency across all optimized target distribution spaces.
+To address this issue, a potential solution is to align the target distributions of all participants’ prior models with that of the prior model producing the most optimal target distribution, referred to as the best prior model.
+Such alignment can be achieved by utilizing an optimizable transformation matrix to map each participant’s target distribution to that of the best prior model (Sun et al., 2024).
+This alignment strategy ensures consistency across all optimized target distributions.
 In summary, it is crucial to first effectively assess each participant’s prior model quality and subsequently train the transformation matrix for alignment.
 A Metric to Quantify Prior Model Quality.
-Drawing inspirations from [32], which proposes an optimizable metric a.k.a.
+Inspired by Wang & Isola (2020), which proposes an optimizable metric a.k.a.
 uniform value loss to achieve feature uniformity on the hypersphere during training, we employ this metric to evaluate the quality of prior models.
-Notably, [32] also provide theoretical validation of the connection between uniformity and feature quality.
+Notably, (Wang & Isola, 2020) also provide theoretical validation of the connection between uniformity and feature quality.
 Specifically, each participant downloads a small shared dataset $S_{X}$ from the platform and computes the uniformity value of their prior model on $S_{X}$.
@@
 $\mathcal{V}_{\text{uniform}}(\boldsymbol{\psi};S)\triangleq\log\mathbb{E}_{\mathbf{x}_{i},\mathbf{x}_{j}\sim S}\left[e^{\tau\|{\boldsymbol{\psi}}(\mathbf{x}_{i})-{\boldsymbol{\psi}}(\mathbf{x}_{j})\|_{2}^{2}}\right],$ (3)
-where $\boldsymbol{\psi}$ is the prior model, $\tau$ is a hyper-parameter set as $2$, consistent with [32].
+where $\boldsymbol{\psi}$ is the prior model, $\tau$ is a hyper-parameter set as $2$, consistent with (Wang & Isola, 2020).
 A lower uniform value indicates a higher-quality prior model, which optimizes targets of superior quality.
-Extensive experiments in Figure 3(c) demonstrate a strong correlation between this metric and the performance of prior, thereby effectively assessing the quality of targets.
+Extensive experiments in Fig.
+5(c) demonstrate a strong correlation between this metric and the performance of prior, thereby effectively assessing the quality of targets.
 Alignment.
 Upon identifying the best prior model, denoted as $\boldsymbol{\psi}^{\star}$, all participants, excluding the best prior model itself, proceed to train an optimizable transformation matrix.
-Specifically, the participant owning the best prior model computes its optimized targets, denoted as ${\mathbf{S}_{Y}}^{\star}$, onthe shared dataset $S_{X}$ using the prior model ${\boldsymbol{\psi}}^{\star}$.
+Specifically, the participant owning $\boldsymbol{\psi}^{\star}$ computes its optimized targets, denoted as $\mathbf{S}_{Y}$$^{\star}$, on the shared dataset $S_{X}$.
 $\mathbf{S}_{Y}$$^{\star}$ are then uploaded to the platform, which ensures they are publicly accessible to all participants.
@@
 where $\mathbf{S}_{X}$ represents the matrix form of $S_{X}$, suitable for input into the network ${\boldsymbol{\psi}}^{(k)}$, and $\mathbf{S}_{Y}$$^{\star}$ also represents the matrix form of $S_{Y}$$^{\star}$.
-After obtaining the transformation matrix $\mathbf{T}^{(k)}$, the participant can convert the optimized targets for its own data using this matrix: ${D_{Y}}^{(k)}=\mathbf{T}^{(k)}\cdot{\boldsymbol{\psi}}^{(k)}({\mathbf{D}_{X}}^{(k)})$, where ${\mathbf{D}_{X}}^{(k)}$ denotes the participant’s subset, and ${D_{Y}}^{(k)}$ are the adjusted targets aligned with the best prior model’s target distribution space.
-As illustrated in Figure 5(b), the proposed alignment strategy effectively mitigates target distribution space inconsistency.
+After obtaining the transformation matrix $\mathbf{T}^{(k)}$, the participant can convert the optimized targets for its own data using this matrix: ${D_{Y}}^{(k)}=\mathbf{T}^{(k)}\cdot{\boldsymbol{\psi}}^{(k)}({\mathbf{D}_{X}}^{(k)})$, where ${\mathbf{D}_{X}}^{(k)}$ denotes the participant’s subset, and ${D_{Y}}^{(k)}$ are the adjusted targets aligned with the best prior model’s target distribution.
+As illustrated in Fig.
+6(b), the proposed alignment strategy effectively mitigates target distribution inconsistency.
 Remark on Privacy.
@@
 Remark on Theoretical Effectiveness.
-Our proposed framework builds upon well-established theoretical foundations.
-Specifically, RELA [29] has theoretically demonstrated that leveraging task-agnostic models, such as pre-trained models, can accelerate learning.
-Notably, our framework diverges substantially from RELA in both its objectives and methodology.
+CoOpt builds upon well-established theoretical foundations.
+In particular, RELA Sun et al.
+(2024) has theoretically demonstrated that leveraging task-agnostic models, such as pre-trained models, can accelerate model learning.
+Nevertheless, CoOpt departs significantly from RELA in both its research objectives and methodological design.
 Specifically, CoOpt introduces a collaborative optimization approach tailored for unlabeled data, targeting a fundamentally distinct problem domain.
 The challenges associated with collaborative data optimization are unique to CoOpt and remain unaddressed by RELA.
-Furthermore, the metric we employ to evaluate the quality of prior models, uniform value loss [32], has been theoretically validated for its effectiveness.
-Table 2: Comparison of CoOpt with Self-Supervised Learning Methods in Accuracy (%) and Training Time (s).
-We use four datasets: CF-10 (CIFAR-10), CF-100 (CIFAR-100), T-IN (Tiny-ImageNet), and IN-1K (ImageNet-1K).
-The best results are marked in bold.
-$\uparrow$ means the performance improvement over the second-best result.
-$\times$ denotes the factor of training speed compared to the second-best result.
-Dataset Metric BYOL DINO MoCo SimCLR SimSiam SwAV DCL CoOpt
-CF-10 Acc. (%) 82.8 $\pm$ 0.1 82.6 $\pm$ 0.0 82.9 $\pm$ 0.1 83.1 $\pm$ 0.0 79.0 $\pm$ 0.0 82.9 $\pm$ 0.1 83.9 $\pm$ 0.1 88.7 $\pm$ 0.1 ($\uparrow$ 4.8)
-Time (s) 1,376.56 1,457.22 1,349.56 1,114.81 1,090.79 1,012.74 1,783.34 540.43 ($\times$ 1.87)
-CF-100 Acc. (%) 51.7 $\pm$ 0.1 51.0 $\pm$ 0.0 57.8 $\pm$ 0.1 55.4 $\pm$ 0.0 44.6 $\pm$ 0.1 53.2 $\pm$ 0.1 58.2 $\pm$ 0.2 65.8 $\pm$ 0.1 ($\uparrow$ 7.6)
-Time (s) 1,406.17 1,419.69 1,425.80 1,103.45 1,139.14 1,072.44 1,701.49 548.11 ($\times$ 1.95)
-T-IN Acc. (%) 43.9 $\pm$ 0.2 36.1 $\pm$ 0.0 42.4 $\pm$ 0.2 41.5 $\pm$ 0.1 40.8 $\pm$ 0.0 39.9 $\pm$ 0.1 44.6 $\pm$ 0.0 58.2 $\pm$ 0.1 ($\uparrow$ 13.6)
-Time (s) 7,086.62 7,030.90 7,133.98 5,621.33 5,531.92 5,540.96 9,201.51 2,852.67 ($\times$ 1.94)
-IN-1K Acc.(%) 61.9 $\pm$ 0.1 52.2 $\pm$ 0.0 57.6 $\pm$ 0.0 58.0 $\pm$ 0.0 55.8 $\pm$ 0.1 57.2 $\pm$ 0.1 60.6 $\pm$ 0.1 68.7 $\pm$ 0.0 ($\uparrow$ 6.8)
-Time (s) 133,766.19 133,156.88 150,420.36 99,176.29 98,656.57 96,134.98 102,450.84 80,096.43 ($\times$ 1.20)
+Furthermore, the metric we employ to evaluate the quality of prior models, uniform value loss (Wang & Isola, 2020), has been theoretically validated for its effectiveness.
 4 Experiments
-In this section, we outline the experimental setup in Section 4.1 and assess the performance of our proposed collaborative data optimization framework, CoOpt, across various real-world scenarios, employing different datasets and architectures, as discussed in Section 4.2.
-Subsequently, we explore a continuous data optimization framework that allows the prior models of participants to evolve, thereby further optimizing the targets, as detailed in Section 4.3.
-Finally, we validate the effectiveness of the introduced uniform value metric and the target alignment strategy in Section 4.4.
+We conduct extensive experiments to demonstrate the key advantages of CoOpt in Sec.
+4.2.
+Specifically, first, to evaluate its efficacy and efficiency in utilizing unlabeled data, we compare CoOpt with state-of-the-art self-supervised learning methods.
+Second, to examine the necessity of distributed optimization, we further compare CoOpt with centralized optimization approaches.
+Third, we train diverse model architectures on the optimized data, thereby assessing its generalizability and reusability.
+Forth, we demonstrate the potential of CoOpt for continuous data optimization, showing how it continuously enhances data quality.
+Furthermore, we explore factors that influence its effectiveness (Sec.
+4.3), including different prior datasets and prior models.
+Finally, we present comprehensive ablation studies (Sec.
+4.4) to verify the impact of each module in CoOpt and its lightweight design.
 4.1 Experimental Setup
 Datasets and Networks:
-We conduct experiments on both large-scale and small-scale datasets, including the full $224\times 224$ ImageNet-1k (224 $\times$ 224) [10], Tiny-ImageNet (64 $\times$ 64) [21], CIFAR-100 [19] and CIFAR-10 (32 $\times$ 32) [20].
-Following previous self-supervised studies [15, 6, 13, 8, 1, 40], we employ a range of model capacities backbone architectures to evaluate the generalizability of our method, including ResNet-18, 50, 101 [16], ViT [12], and a series of CLIP-based models [26].
+We conduct experiments on both large-scale and small-scale datasets, including ImageNet-1k (224 $\times$ 224) (Deng et al., 2009), Tiny-ImageNet (64 $\times$ 64) (Le & Yang, 2015), CIFAR-100 (Krizhevsky et al., 2009a) and CIFAR-10 (32 $\times$ 32) (Krizhevsky et al., 2009b).
+Following previous self-supervised studies (He et al., 2020; Chen et al., 2020a; Grill et al., 2020; Chen & He, 2021; Assran et al., 2023; Zhang et al., 2024), we employ a range of backbone architectures to evaluate the generalizability of our method, including ResNet-{18, 50, 101} (He et al., 2016), ViT (Dosovitskiy et al., 2020), and a series of CLIP-based models (Radford et al., 2021).
 Baselines:
-For the unlabeled data, referring to a prior widely-used benchmark [9], we consider several state-of-the-art self-supervised methods as baselines for a broader practical impact, including: SimCLR [6], BYOL [13], DINO [4], MoCo [15], SimSiam [8], SwAV [3], and DCL [36].
-Notably, we do not compare our approach with knowledge distillation (KD) or dataset distillation (DD) methods, as summarized in Table 1, since they primarily foucus on the optimization of labeled data.
-In contrast, this work focuses on the utilization of unlabeled data.
+For the unlabeled data, following a widely used benchmark (Da Costa et al., 2022), we compare against state-of-the-art self-supervised (SSL) methods, including: SimCLR (Chen et al., 2020a), BYOL (Grill et al., 2020), DINO (Caron et al., 2021), MoCo (He et al., 2020), SimSiam (Chen & He, 2021), SwAV (Caron et al., 2020), and DCL (Yeh et al., 2022).
+Notably, we do not compare with knowledge distillation (KD) or dataset distillation (DD) methods, since the training paradigm of KD differs significantly from ours, while DD primarily focuses on the labeled data.
 Evaluation and Metrics:
-Following previous benchmarks [15, 6, 13, 8], we evaluate the representation quality of the trained models by evaluating their test accuracy (%) using an offline linear probing strategy.
+Following previous benchmarks (Grill et al., 2020; Chen & He, 2021), we evaluate the representation quality of models by evaluating their test accuracy (%) using an offline linear probing strategy.
 Additionally, computational efficiency quantified by time cost (s).
@@
 The proposed algorithm, CoOpt, involves an open data platform and facilitates interaction among multiple participants.
-The implementation follows five key steps (detailed in Section 3.2), and more details are provided in Appendix B.
+The implementation follows five key steps (detailed in Sec.
+3.2), and more details are provided in App.
+B.
   1. 1.
@@
     For steps 2 and 3, in practical applications, each participant can use publicly pre-trained models or their own models directly as the prior model.
-To simulate the diversity of prior models in practical applications, we utilize pre-trained models trained on ImageNet as well as a series of pre-trained CLIP-based models.
+To simulate the diversity of prior models in practical applications, we use a series of pre-trained CLIP-based models.
   3. 3.
@@
 For fair comparisons, all methods in the experiments are executed with the same hyperparameters.
-4.2 Comparison with the state-of-the-art methods
-We evaluate our framework, CoOpt, across various scenarios: (1) participants leverage a diverse array of pre-trained prior models, characterized by variations in datasets, architectures, and training paradigms.
-(2) we specifically evaluate the robustness of our method when employing different prior datasets used for training prior models; (3) finally, we explore special cases involving human or weak models, particularly where there are resource-rich or resource-poor participants.
-Diverse Pre-trained Prior Models.
-Given the diversity of prior models, we employ a comprehensive set of pre-trained models, including those trained on ImageNet-1K and CLIP-based models.
-Details of these models are provided in Appendix B.2.
-These models are trained using various training paradigms, datasets, and architectures.
-As shown in Table 2, our method CoOpt demonstrates superior performance and training efficiency compared to existing self-supervised learning methods.
-Additional experimental results for the optimized data, trained across a wider range of architectures, are presented in Table 5 within Appendix B.3.
-  1. (a)
-    Performance: CoOpt significantly outperforms state-of-the-art self-supervised learning methods.
-For instance, it achieves an improvement of 13.6% and 6.8% over the leading self-supervised approach BYOL on Tiny-ImageNet and ImageNet-1K, respectively.
-Notably, CoOpt demonstrates more significant improvements on larger-scale datasets, such as Tiny-ImageNet and ImageNet-1K, which is particularly advantageous given the current emphasis on large-scale data era.
-  2. (b)
-    Efficiency: CoOpt demonstrates a substantial reduction in training costs across various datasets.
-Notably, on the Tiny-ImageNet, CoOpt achieves a training speed that surpasses BYOL by a factor of approximately ×$\times 1.94$.
+4.2 What are the advantages of CoOpt?
+Table 1: Comparison of CoOpt with Self-Supervised Learning Methods in Accuracy (%) and Training Time (s).
+We use four datasets: CF-10 (CIFAR-10), CF-100 (CIFAR-100), T-IN (Tiny-ImageNet), and IN-1K (ImageNet-1K).
+The best results are marked in bold.
+$\uparrow$ means the performance improvement over the second-best result.
+$\times$ denotes the factor of training speed compared to the second-best result.
+Dataset Metric BYOL DINO MoCo SimCLR SimSiam SwAV DCL CoOpt (Ours)
+CF-10 Acc. (%) 82.8 $\pm$ 0.1 82.6 $\pm$ 0.0 82.9 $\pm$ 0.1 83.1 $\pm$ 0.0 79.0 $\pm$ 0.0 82.9 $\pm$ 0.1 83.9 $\pm$ 0.1 89.5 $\pm$ 0.1 ($\uparrow$ 5.6)
+Time (s) 1,376.56 1,457.22 1,349.56 1,114.81 1,090.79 1,012.74 1,783.34 540.43 ($\times$ 1.87)
+CF-100 Acc. (%) 51.7 $\pm$ 0.1 51.0 $\pm$ 0.0 57.8 $\pm$ 0.1 55.4 $\pm$ 0.0 44.6 $\pm$ 0.1 53.2 $\pm$ 0.1 58.2 $\pm$ 0.2 67.3 $\pm$ 0.1 ($\uparrow$ 9.1)
+Time (s) 1,406.17 1,419.69 1,425.80 1,103.45 1,139.14 1,072.44 1,701.49 548.11 ($\times$ 1.95)
+T-IN Acc. (%) 43.9 $\pm$ 0.2 36.1 $\pm$ 0.0 42.4 $\pm$ 0.2 41.5 $\pm$ 0.1 40.8 $\pm$ 0.0 39.9 $\pm$ 0.1 44.6 $\pm$ 0.0 60.3 $\pm$ 0.1 ($\uparrow$ 15.7)
+Time (s) 7,086.62 7,030.90 7,133.98 5,621.33 5,531.92 5,540.96 9,201.51 2,852.67 ($\times$ 1.94)
+IN-1K Acc.(%) 61.9 $\pm$ 0.1 52.2 $\pm$ 0.0 57.6 $\pm$ 0.0 58.0 $\pm$ 0.0 55.8 $\pm$ 0.1 57.2 $\pm$ 0.1 60.6 $\pm$ 0.1 69.8 $\pm$ 0.1 ($\uparrow$ 7.9)
+Time (s) 133,766.19 133,156.88 150,420.36 99,176.29 98,656.57 96,134.98 102,450.84 80,096.43 ($\times$ 1.20)
+Comparison with SSL Methods.
+As shown in Tab.
+1, our CoOpt demonstrates superior performance and efficiency compared to existing self-supervised learning methods.
+We also visualize the training dynamic in Fig.
+5(a).
+Specifically, CoOpt achieves an improvement of 7.9% over the leading self-supervised approach BYOL on ImageNet-1K.
+For Efficiency, CoOpt demonstrates a substantial improvememt across various datasets.
+Notably, on the Tiny-ImageNet, CoOpt achieves a training speed that surpasses the efficient method SimSiam by a factor of approximately $\times 1.94$.
+Table 2: Comparison with Centralized Optimization.
+Method Time (s) Acc. (%)
+Centralized 23.71 62.1 $\pm$ 0.1
+Ours 16.31 65.8 $\pm$ 0.1
+Comparison with Centralized Optimization.
+A key advantage of CoOpt lies in its ability to use diverse prior models in parallel, thereby enabling efficient optimization.
+To validate this, we compare CoOpt with centralized optimization, where a single model is used to optimize all unlabeled data.
+Specifically, we consider 10 prior models of varying quality on CIFAR-100.
+For the centralized setting, we report the mean performance of these 10 independently selected models to ensure fairness.
+As summarized in Tab.
+2, CoOpt consistently demonstrates superior efficiency and efficacy, verifying the benefits of distributed over centralized optimization.
+This improvement arises from the proposed target alignment strategy (Sec.
+3.4), which leverages high-quality priors to enhance the target distribution of weaker models.
+While one might envision an ideal centralized solution that exclusively employs the best prior model, such an approach is rarely practical in real-world scenarios A major concern is fairness and computational burden, as concentrating all computation on a single party imposes excessive cost and discourages participation.
+Another challenge is privacy, since high-performing models are typically proprietary, and centralizing them may violate ownership or data-sharing constraints.
+In contrast, CoOpt collaboratively leverage diverse prior models, achieving competitive performance while substantially reducing the burden on any individual participant.
+Table 3: Comparison of CoOpt with BYOL on Diverse Networks.
+Method
+Network BYOL CoOpt
+ResNet-50 60.4 $\pm$ 0.1 63.8 $\pm$ 0.0
+ResNet-101 61.5 $\pm$ 0.2 65.7 $\pm$ 0.2
+MobileNet-v2 24.0 $\pm$ 0.5 58.1 $\pm$ 0.0
+Efficientnet-b0 2.3 $\pm$ 0.2 70.7 $\pm$ 0.2
+ViT 38.5 $\pm$ 0.1 57.8 $\pm$ 0.1
+Generalizability and Reusability of Optimized Data.
+Another key advantage of our optimized data lies in its strong generalizability and reusability: once constructed, it can be directly employed for downstream diverse model training without further modification.
+To evaluate this advantage, we conduct experiments by training a variety of neural architectures on the optimized data and compare the results to a strong baseline, BYOL, which relies on training from scratch on the original unlabeled dataset.
+The results are summarized in Tab.
+3.
+Obviously, CoOpt consistently delivers significant performance improvements over BYOL across multiple architectures.
+In particular, BYOL suffers substantial degradation when applied to lightweight networks such as MobileNet-v2.
+A plausible explanation is that these models are more sensitive to unstable batch normalization (BN) statistics in early network layers.
+Instead, our optimized data exhibits strong generalization across diverse architectures.
+Figure 4: A Practical Scenario: Continuous Optimization.
+Continuous Data Optimization.
+We further explore a practical scenario where the prior models undergo temporal evolution.
+For example, a participant’s initial model, such as ResNet-50, might be upgraded to a higher-capacity model like ResNet-101 as their computational resources improve.
+Consequently, the process can be treated as a dynamic and continuous procedure.
+The detailed description of the process is provided in App.
+B.4.
+We simulate this scenario by making random 2020% of the participants increase their model capacity in each round.
+The training curves across 10 rounds on CIFAR-100 are shown in Fig.
+4.
+The results demonstrate that in CoOpt, as the prior models evolve, the quality of the targets improves, thereby facilitating continuous optimization.
+In particular, over 10 rounds, the continuous optimization setting yields a 4.6% performance gain.
+4.3 What influence the effectiveness of CoOpt?
+We investigate the key factors that influence the effectiveness of the optimized data.
+Since the optimization of unlabeled data in CoOpt relies solely on task-agnostic prior models, we focus on two primary aspects in these experiments: prior datasets and prior models.
+(a) Training Curves
+(b) Varying Prior Models
+(c) Uniform Value
+(d) Shared Data Size
+Figure 5: Comprehensive Analysis of CoOpt.
+(a) Training curves: Comparison of SSL methods and our CoOpt.
+(b) Prior Models With Varying Accuracies.
+Even with a very weak prior model, CoOpt accelerates the early-stage training.
+(c) Correlation Verification: Verify the correlation between the uniform value and performance.
+(d) Influence of shared data size: As shared data’s size increases, the performance gains diminish.
 Table 4: Comparison of CoOpt with BYOL Across Diverse Prior Datasets.
@@
 All models are based on ResNet-18.
-Dataset BYOL (Baseline) Our CoOpt (Diverse Prior Datasets)
+Dataset
+BYOL
+Our CoOpt (Diverse Prior Datasets)
 CIFAR-10 (P) CIFAR-100 (P) Tiny-ImageNet (P) ImageNet-1K (P)
@@
 ImageNet-1K 61.9 $\pm$ 0.1 31.7 $\pm$ 0.1 ($\downarrow 30.2$) 31.8 $\pm$ 0.0 ($\downarrow 30.1$) 40.5 $\pm$ 0.0 ($\downarrow 21.4$) 71.2 $\pm$ 0.0 ($\uparrow$ 9.3)
-Diverse Prior Datasets.
-In the last experiment, we employ pre-trained models, some of which are congruent with the unlabeled training data, while others are incongruent.
-We refer to these datasets as the prior datasets.
-To rigorously evaluate the influence of these prior datasets, we perform an analysis across scenarios where the prior datasets used for prior models either align with or differ from the unlabeled training dataset.
+Influence of Prior Datasets.
+To rigorously evaluate the influence of prior datasets that are used for training prior models, we perform an analysis across scenarios where the prior datasets used for prior models either align with or differ from the unlabeled training dataset.
 For instance, in the aligned scenario, the training dataset is CIFAR-10, and the prior models are also trained on CIFAR-10 (CIFAR-10 (P)).
 Conversely, in the divergent scenario, the training dataset remains CIFAR-10, while the prior models are trained on CIFAR-100 (P).
-Details of these models are provided in Appendix B.4.
-We evaluate our approach on four publicly available datasets, with the results summarized in Table 3.
-  1. (a)
-    Utilizing prior models trained on the same dataset as the target training data can lead to notable performance gains.
-  2. (b)
-    For complex training datasets, leveraging prior models trained on simpler datasets may degrade performance compared to BYOL This degradation arises due to the limited informativeness of simpler prior datasets, which provide weaker guidance.
-  3. (c)
-    Importantly, for all unlabeled training datasets, employing prior models trained on ImageNet-1K consistently yields significant performance gains.
-This can be attributed to the strong generalization abilities of such models.
-This observation is particularly pertinent in practical applications, as the majority of publicly available pre-trained models are derived from ImageNet-1K.
+Details of these models are provided in App.
+B.3.
+We evaluate our approach on four publicly available datasets, with the results summarized in Tab.
+4.
+Obviously, for all unlabeled training datasets, employing prior models trained on ImageNet-1K consistently yields notable performance gains, attributed to their strong generalization abilities.
+This observation is particularly relevant in practical applications, as most publicly available pre-trained models are derived from ImageNet-1K or even larger datasets.
+On the other hand, for complex training datasets, leveraging prior models trained on simpler datasets may result in degraded performance compared to BYOL.
+This is likely due to the limited informativeness of simpler prior datasets, which provide weaker guidance.
+We further examine the influence of weak models in Fig.
+5(b).
 Table 5: Comparison of CoOpt with BYOL in Presence of Human or Weak Prior Models.
@@
 BYOL 82.8 $\pm$ 0.1
-CoOpt 88.7 $\pm$ 0.1
-89.0 $\pm$ 0.0
-90.3 $\pm$ 0.0
-89.2 $\pm$ 0.0
-Extreme Cases of Prior Models: Human or Weak.
+CoOpt 89.5 $\pm$ 0.1
+89.2 $\pm$ 0.2
+90.5 $\pm$ 0.1
+89.8 $\pm$ 0.1
+Special Cases of Prior Models: Human or Weak Involvement.
 In real-world applications, extreme cases arise due to the varying capabilities of participants.
-For example, some participants have extensive resources and employ human annotators for labeling, while others may have limited resources and rely on weak models with inferior performance and generalization abilities.
-In our experiments, we define weak models as those trained during intermediate stages that are even far from convergence.
+For example, some participants have extensive resources and can employ human annotators for labeling, while others may have limited resources and rely on weak models with inferior generalization abilities.
+In this experiment, we define weak models as those trained during intermediate stages that are even far from convergence.
 To simulate the conditions, in addition to the prior models used in the first experiment, we incorporate $5$ prior models, either human or weak models to the data optimization process.
-The results are summarized in Table 4.
+The results are summarized in Tab.
+5.
 Surprisingly, even the inclusion of weaker models contributes to enhancing the final performance, indicating that such models can still provide valuable information.
 Moreover, it is important to note that the integration of high-capacity, human-like models results in significant performance improvements.
-(a) Training curves
-(b) Varying Prior Models
-(c) Uniform value
-(d) Shared data size
-Figure 3: Comprehensive Analysis of CoOpt.
-(a) Training curves: Comparison of various self-supervised learning methods and the proposed CoOpt.
-(b) Prior Models With Varying Accuracies.
-Employing a strong model as the prior model is not essential (c) Correlation Verification: Examine the relationship between the uniform value and performance, which demonstrates a strong negative correlation.
-(d) Influence of shared data size: As the size of the shared data increases, the performance gains diminish.
-Moreover, we conduct experiments with only weaker model, as shown in Figure 3(b).
-Here, The dashed line represents the training curve of BYOL (baseline), while the solid lines correspond to prior models with different accuracies.
-"Prior model (BYOL)" indicates the use of BYOL as the prior model, and the accuracies of the other prior models are all lower than that of BYOL.
-Notably, most prior models surpass BYOL performance in the early stages of training, and even a model with approximately 75% accuracy can outperform BYOL in the final performance.
-The results demonstrate that employing a strong model as the prior model is not essential.
-Figure 4: A Practical Scenario: Continuous Optimization Prior models experience temporal evolution, leading to improved target quality across multiple rounds.
-4.3 Continuous Data Optimization
-We explore another practical scenario where the prior models undergo temporal evolution.
-For example, a participant’s initial model, such as ResNet-50, might be upgraded to a higher-capacity model like ResNet-101 as their computational resources improve.
-Consequently, the process can be treated as a dynamic and continuous procedure.
-The detailed description of the process is provided in Appendix B.5.
-We simulate this scenario by conducting experiments where, in each interaction round between the platform and participants, $20$% of the participants randomly update their model architectures to reflect an increase in model capacity.
-The training curves across 10 rounds on CIFAR-100 are depicted in Figure 4.
-The results have demonstrated that in CoOpt, as the prior models evolve, the quality of the targets improves, thereby facilitating continuous optimization.
-In particular, compared to the baseline BYOL, these improvements can achieve an enhancement of $4.6$%.
-4.4 Ablation Study
-Effectiveness of Uniform Value.
-To evaluate the effectiveness of uniform value in estimating the quality of prior models, we employ a diverse set of prior models, calculating both their uniform value and test accuracy.
-We employ the Spearman rank correlation coefficient [^4] The formula is: $\rho=1-\frac{{6\sum d_{i}^{2}}}{{n(n^{2}-1)}}$, where $d_{i}$ represents the difference between the ranks of each pair of observations and $n$ denotes the number of observations.
-$\rho$ [39] to quantify the association between Uniform Value and accuracy.
-As illustrated in Figure 3(c), the Spearman rank correlation coefficient is $\rho=-0.9714$, indicating a strong correlation between uniform value and model performance, thereby effectively assessing the quality of prior models.
-Target Distribution Alignment.
-In the real-world applications, participants often employ diverse prior models, which results in target distribution space inconsistencies, as demonstrated in Figure 5(a).
-We conduct an ablation study w/o alignment and w/ alignment to demonstrate the importance and necessarily of target distribution space alignment, and the training curves presented in Figure 5(d).
-Three key observations include:
-  1. (a)
-    Compared to without alignment (green line), our proposed approach (blue line) yields a performance improvement of $16.9$%, which highlights the critical importance of aligning the target distribution space.
-Notably, the training curve without alignment (green line) initially ascends before declining, suggesting that during the early phase of model training, optimizing the target leads to performance enhancement.
-However, as training continues, the severely inconsistent targets significantly degrade model performance.
-  2. (b)
-    Furthermore, we compare our alignment strategy, which aligns to the best prior model, with two other straightforward strategies: aligning to a medium prior model (purple line) and aligning to the worse prior model (red line).
-Obviously, three strategies outperform the scenario without alignment, and aligning to the best prior model provides the most significant performance gains.
-  3. (c)
-    To further analyze the underlying reasons, we employ t-SNE visualization.
-Notably, a comparison between alignment with the better participant ( Figure 5(b)) and the worse participant ( Figure 5(c)) reveals that alignment with a worse-quality model diminishes the representative capability of the targets.
-This, in turn, results in less effective guidance for model training.
-Influence of Shared Unlabeled Data Size.
-The shared unlabeled data $S$ is used to estimate the uniform value and compute the transformation matrix for target alignment, as detailed in Section 3.4.
-To explore the influence of the size of shared unlabeled data, we conduct experiments on both CIFAR-10 and CIFAR-100, varying the size of the shared data in proportions of $\{0.01,0.05,0.1,0.2,0.4,0.6,0.8\}$ relative to the original dataset.
-The results are presented in Figure 3(d).
-It is observed that as the size of the shared data increases, performance gains become marginal.
-Specifically, increasing the proportion of data from $0.01$% to $0.05$% results in improvements of $2.4$% and $2.5$% for CIFAR-10 and CIFAR-100, respectively.
-However, increasing the proportion from $0.1\%$ to $0.8\%$ results in only a $0.5\%$ and $0.6\%$ improvement.
-This suggests that a small-sized shared unlabeled dataset is adequate for accurate uniform value estimation and the computation of the transformation matrix.
 (a) Before Alignment
@@
 (d): Training curves with and without alignment.
-5 Conclusion and Future Work
-Conclusion.
-In this paper, we introduce CoOpt, a pioneering data-centric, efficient, and parallelized framework for collaborative optimization of unlabeled data.
-This data-centric approach results in architecture-agnostic optimized data that are reusable across diverse network architectures, while simultaneously reducing the number of training iterations required, thereby enhancing overall computational efficiency.
-The collaborative mechanism effectively exploits distributed computation to substantially lower the cost per participant.
+Extreme Cases of Prior Models: All are Weak.
+Moreover, we conduct experiments with only weak model, as shown in Fig.
+5(b).
+Here, the dashed line represents the training curve of BYOL (baseline), and the solid lines means prior models with different accuracies.
+"Prior model (BYOL)" indicates the use of BYOL as the prior model, and the accuracies of the other prior models are all lower than that of BYOL.
+While a stronger prior model does yield better performance, our results demonstrate that even with a moderately weak prior model with approximately 75% accuracy, our method can still outperform the baseline.
+More importantly, even when prior models are substantially weak, CoOpt still significantly outperform baseline in the early training stages.
+4.4 Ablation Study
+Can Uniform Value Effectively Assess the Quality of Prior Models?
+To evaluate the effectiveness of uniform value in estimating the quality of prior models, we employ a diverse set of prior models and compute both their uniform values and corresponding test accuracies.
+To quantify the relationship between these two measures, we adopt the Spearman rank correlation coefficient [^3] The formula is: $\rho=1-\frac{{6\sum d_{i}^{2}}}{{n(n^{2}-1)}}$, where $d_{i}$ represents the difference between the ranks of each pair of observations and $n$ denotes the number of observations.
+$\rho$ (Zar, 2014) to quantify the association between Uniform Value and accuracy.
+As illustrated in Fig.
+5(c), the Spearman rank correlation coefficient is $\rho=-0.9714$, indicating a strong correlation between uniform value and model performance, thereby effectively assessing the quality of prior models.
+Why Target Distribution Alignment is Necessary?
+In real-world applications, participants often adopt diverse prior models, leading to inconsistencies in the target space (Fig.
+6(a)).
+To verify the necessity of alignment, we conduct an ablation study with and without alignment, and the training curves in Fig.
+6(d) show that our alignment method improves performance by 16.9%, veifying its effectiveness.
+Furthermore, We compare three strategies: aligning to the best prior model (ours), a medium-quality prior model (purple line), and the worst prior model (red line).
+All strategies outperform the unaligned baseline, but aligning to the best prior model yields the largest performance gains, as reported in Tab.
+7 of App.
+B.5.
+To analyze the underlying reasons, we employ t-SNE visualization.
+As shown in Fig.
+6(b) and Fig.
+6(c), alignment with a high-quality model enhances the representational capability of the targets, whereas alignment with a poor model diminishes it.
+How Much Shared Unlabeled Data Is Enough?
+The shared unlabeled data $S$ is used to estimate the uniform value and compute the transformation matrix for target alignment, as detailed in Sec.
+3.4.
+To explore the influence of the data size, we vary the proportion of shared data on CIFAR-10 and CIFAR-100, as shown in Fig.
+5(d).
+Obviously, as the size of the shared data increases, performance gains become marginal, indicating that a very small fraction (around 0.05%) is sufficient for accurate estimation and alignment.
+We further validate this on ImageNet-1K (Tab.
+8 in App.
+B.6), where only 0.001% of the data achieves comparable results to larger proportions.
+This minimal requirement imposes negligible additional overhead, ensuring scalability to very large datasets.
+Table 6: Comparison of Time (s) on ImageNet-1K.
+BYOL Uniform value Alignment
+133,766.19 139.16 36.97
+Does Alignment Incur Significant Overhead?
+We report the computational cost of uniform value estimation and alignment in Tab.
+6, verifying that both incur only negligible cost.
+This efficiency stems from the former requiring just a single forward pass to obtain targets, while the latter involves optimizing a lightweight matrix.
+5 Conclusion
+We introduce CoOpt, a pioneering data-centric, parallelized, and efficient framework for collaborative optimization of unlabeled data.
+This data-centric approach results in architecture-agnostic optimized data that are reusable across diverse network architectures, while simultaneously reducing the number of training iterations required, thereby enhancing overall efficiency.
 Furthemore, within CoOpt, we identify a critical issue: Target Distribution Inconsistency, which arises from the diversity of prior models used in data optimization.
-To mitigate this, we propose a target alignment strategy.
-Extensive experiments conducted across various real-world scenarios demonstrate the superior effectiveness and efficiency of the CoOpt framework across diverse datasets and architectures.
-One limitation is that the experiments are primarily conducted on image datasets, without extending to other domain data.
-Future Work.
-In the future, our objectives include the following: (1) Develop advanced methodologies to ensure strong privacy guarantees; (2) Conduct an in-depth exploration of continuous data optimization through iterative processes, enabling dynamic collaboration among participants.
-(3) Integrate diverse domain datasets into the optimization framework.
-(3) Explore more practical scenarios, such as participants optimizing their local data and then uploading it to the shared platform.
-This paradigm can broaden the platform’s functionalities, where organizations and individuals can contribute to and benefit from a globally data ecosystem.
+To mitigate this, we propose a lightweight target alignment strategy.
+Extensive experiments demonstrate the superior effectiveness and efficiency of the CoOpt framework across diverse datasets and architectures.
+One limitation is that when all prior models are extremely weak, the overall performance inevitably degrades.
+As future work, we aim to develop advanced strategies to more effectively exploit optimized data derived from all extremely weak priors, with our current study verifying efficiency in the early training stage.
 Appendix A Related Work
 Self-supervised Learning: A Model-Centric Perspective.
-Self-supervised learning [7] aims to exploit the intrinsic relationships within unlabeled data.
-For example, InstDisc [34] uses instance discrimination as a pretext task.
-Building on this, CMC [30] proposes to use multiple views of an image as positive samples and take another one as the negative.
-MoCo [15] significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples.
-SimCLR [6] highlights the importance of hard positive sample strategies by introducing data augmentation.
-Notably, BYOL [13] discards negative sampling and surpasses the performance of SimCLR [6].
+Self-supervised learning (Chen et al., 2020b) aims to exploit the intrinsic relationships within unlabeled data.
+For example, InstDisc (Wu et al., 2018) uses instance discrimination as a pretext task.
+Building on this, CMC (Tian et al., 2020) proposes to use multiple views of an image as positive samples and take another one as the negative.
+MoCo (He et al., 2020) significantly increases the number of negative samples but uses a simplistic strategy for selecting positive samples.
+SimCLR (Chen et al., 2020a) highlights the importance of hard positive sample strategies by introducing data augmentation.
+Notably, BYOL (Grill et al., 2020) discards negative sampling and surpasses the performance of SimCLR (Chen et al., 2020a).
 Summary. They are constrained to specific architectures and incur high computational costs due to the reliance on large batch sizes or memory banks.
 Knowledge Distillation: Optimzing Targets.
-Knowledge distillation [17] employs soft labels generated by teacher models to improve the performance of a student model and expedite its training [11].
+Knowledge distillation (Hinton, 2015) employs soft labels generated by teacher models to improve the performance of a student model and expedite its training (Dong et al., 2023).
 Many following works aim to enhance the use of soft labels for more effective knowledge transfer.
-For example, WSLD [45] analyzes soft labels and distributes different weights for them from a perspective of bias-variance trade-off.
-DKD [44] decouples the logits and assigns different weights for the target and non-target classes.
-Moreover, several studies [37, 11] demonstrate that knowledge distillation can accelerate the training process.
+For example, WSLD (Zhou et al., 2021) analyzes soft labels and distributes different weights for them from a perspective of bias-variance trade-off.
+DKD (Zhao et al., 2022) decouples the logits and assigns different weights for the target and non-target classes.
+Moreover, several studies (Yim et al., 2017; Dong et al., 2023) demonstrate that knowledge distillation can accelerate the training process.
 Dataset Distillation: Optimizing Both Samples and Targets.
-Dataset distillation [33] aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost.
-The majority of methods focus on optimizing images, which can be categorized into five primary approaches [22]: meta-learning frameworks [33, 46], gradient matching [43, 41], distribution matching [42, 38], trajectory matching [5, 14], and decoupling frameworks [38, 29].
-Recently, some studies [27, 25] have shifted focus towards label distillation, aiming to obtain high-quality soft labels.
+Dataset distillation (Wang et al., 2018) aims to learn a compact distilled dataset that can achieve comparable performance to the original dataset with less training cost.
+The majority of methods focus on optimizing images, which can be categorized into five primary approaches (Lei & Tao, 2023): meta-learning frameworks (Wang et al., 2018; Zhou et al., 2022), gradient matching (Zhao et al., 2020; Zhao & Bilen, 2021), distribution matching (Zhao & Bilen, 2023; Yin et al., 2023), trajectory matching (Cazenavette et al., 2022; Guo et al., 2024), and decoupling frameworks (Yin et al., 2023; Sun et al., 2024).
+Recently, some studies (Shang et al., 2025; Qin et al., 2024) have shifted focus towards label distillation, aiming to obtain high-quality soft labels.
 This approach has demonstrated notable efficiency and effectiveness.
-Summary. Both approaches are computationally expensive as they require task-specific models.
 Appendix B Experimental Details
@@
 B.2 Diverse Pre-trained Prior Models.
-We utilize 10 pre-trained prior models, which are downloaded using the torchvision package.
-These models consist of 5 ImageNet-1K-based models, varying from small-scale to large-scale architectures, and 5 CLIP-based model.
+We utilize 10 CLIP-based pre-trained prior models varying from small-scale to large-scale architectures, which are downloaded using the torchvision package.
 Unlabeled data is equally split among participants.
-B.3 Diverse Network for Model training in Optimized data.
-Table 5: Comparison of CoOpt with BYOL for Diverse Network on CIFAR-100.
-Method
-Network BYOL CoOpt
-ResNet-50 60.4 $\pm$ 0.1 63.8 $\pm$ 0.0
-ResNet-101 61.5 $\pm$ 0.2 65.7 $\pm$ 0.2
-MobileNet-v2 24.0 $\pm$ 0.5 58.1 $\pm$ 0.0
-Efficientnet-b0 2.3 $\pm$ 0.2 70.7 $\pm$ 0.2
-ViT 38.5 $\pm$ 0.1 57.8 $\pm$ 0.1
-We conduct experiments across various network architectures, comparing our approach, CoOpt, with BYOL, as presented in Table 5.
-Notably, CoOpt consistently demonstrates significant performance improvements over BYOL across multiple architectures.
-Furthermore, BYOL exhibits markedly inferior performance when applied to lightweight architectures such as MobileNet-v2 and EfficientNet-b0, in contrast to its performance with larger backbones like ResNet-50.
-One plausible reason for this performance drop is the instability of batch normalization (BN) statistics in the early layers of these lightweight networks.
-Specifically, the narrow layers (i.e., those with small channel width) in early stages produce activations with limited sample size per channel, resulting in high variance in the estimated batch mean and variance.
-Since BYOL relies on two differently augmented views of the same image, this BN-induced instability leads to exaggerated discrepancies between the two projected features, which deviates from the intended invariance objective of the method.
-This mismatch potentially impairs the optimization process and harms representation quality.
-B.4 Diverse Prior Datasets
+B.3 Diverse Prior Datasets
 In this set of experiments, we investigate the effectiveness of our proposed CoOpt when the prior datasets are either similar to or distinct from the training datasets.
@@
 The results are then aggregated on the open platform.
-B.5 Continuous Data Optimization
+B.4 Continuous Data Optimization
 Figure 7: Common Target Conflict Scenarios
@@
 Iterative Optimization Process
-The overall process is illustrated in Figure 6. The framework operates as a loop:
+The overall process is illustrated in Fig. 7. The framework operates as a loop:
   1. 1.
@@
     This process repeats over multiple interaction rounds, progressively enhancing the dataset.
+Table 7: Performance Under Three Selection Scenarios.
+Dataset BYOL No align Best Medium Worse
+CIFAR-100 51.7 $\pm$ 0.1 44.7 $\pm$ 0.0 65.3 $\pm$ 0.0 63.1 $\pm$ 0.0 60.1 $\pm$ 0.1
+B.5 Various Alignment strageties
+As shown in Fig.
+6(d), even when the best prior model is mis-selected (i.e., prior models are aligned to a medium-performing prior or even the worst prior, represented by the purple and red lines, respectively), the overall performance still surpasses the “no align” baseline.
+Furthermore, we present the final performance in Tab.
+7.
+The results suggest that our method continues to outperform the baseline methods under three selection scenarios and demonstrates robustness to prior model mis-selection.
+Table 8: Influence of Shared Data Size on Large-Scale ImageNet-1K.
+Dataset BYOL 0.001 0.1
+ImageNet-1K 61.9 $\pm$ 0.1 68.8 $\pm$ 0.1 68.7 $\pm$ 0.0
+B.6 Size of Shared Data
+The shared unlabeled data $S$ is used to estimate the uniform value and compute the transformation matrix for target alignment, as detailed in Sec.
+3.4.
+The shared dataset is public and randomly selected.
+This eliminates privacy concerns and does not require matching any specific data distribution, thus avoiding issues of distribution mismatch or privacy leakage.
+Moreover, we further evaluated our method on the large-scale ImageNet-1K.
+As shown in Tab.
+8, only 0.001% of the data samples are sufficient to achieve performance comparable to using larger amounts of data.
+This minimal requirement imposes negligible additional overhead, ensuring scalability to very large datasets.